### v1

Abstract
For the numerical solution of nonsmooth problems, sometimes it is not necessary that an exact subgradient/generalized Jacobian is at our disposal, but that a certain semismoothness property is fulfilled. In this paper we consider not only semismoothness of nonsmooth real- and vector-valued mappings, but also its interplay with the semismoothness property for multifunctions. In particular, we are interested in the semismoothness of solution maps to parametric semismooth inclusions. Our results are expressed in terms of suitable generalized derivatives of the set-valued part, i.e., by limiting coderivatives or by SC (subspace containing) derivatives. As a byproduct we identify a class of multifunctions having the remarkable property that they are strictly proto-differentiable almost everywhere (with respect to some Hausdorff measure) on their graph.
Key words. Nonsmooth problems, semismoothness, solution map, generalized derivative
MSC codes. 65K10, 65K15, 90C26.
1 Introduction
For solving problems of the form
$\min\theta(x)\quad\mbox{ subject to }\quad x\in U_{\rm ad},$
where $\theta:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a locally Lipschitzian function and $U_{\rm ad}\subset\mathbb{R}^{n}$ is a closed set, the use of bundle methods is a popular approach. There exists a bunch of bundle methods in the literature, cf. [17, 14, 29], and all these methods have in common that they require an
oracle, which returns for arbitrary $x\in\mathbb{R}^{n}$ the function value $\theta(x)$ and one subgradient $g(x)\in\partial_{c}\theta(x)$, (1)
where $\partial_{c}\theta$ denotes the Clarke subdifferential. Then, under the additional assumption that
$\theta$ is weakly semismooth,
it can be shown that the bundle method either stops after finitely many steps at a C-stationary point $\bar{x}$ or, if the produced sequence $x_{k}$ remains bounded, one of its accumulation point is C-stationary.
It seems to be that Dempe and Bard [4] were the first who observed that the Bundle-Trust region algorithm (BT algorithm) of Schramm and Zowe [29] also works, when the oracle only delivers a so-called pseudosubgradient. Pseudosubgradients were introduced by Norkin [22, 23] and have the remarkable feature that they fulfill a certain semismoothness property. Although this property does not necessarily imply that $\theta$ is weakly semismooth, one can show that a similar convergence result holds as above: If the BT-algorithm produces a bounded sequence then at least one accumulation point is pseudo-stationary, i.e., 00 is a pseudosubgradient of $\theta$ at this point. By checking the convergence proof for other bundle methods [14, 17] one can show that a similar assertion hold: If the oracle delivers only a pseudosubgradient, one can still prove that the method works and returns pseudo-stationary points. This is important for applications: Whereas the calculus for the Clarke subdifferential often yields only inclusions, the calculus for pseudosubdifferentials yields again a pseudosubdifferential.
Semismoothness also plays an important role in another field of numerical analysis, namely the solution of nonsmooth equations
$G(x)=0,$
where $G:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a locally Lipschitz mapping. Whereas in the original semismooth Newton method introduced by Qi and Sun [26] matrices from the Clarke generalized Jacobian were used, it was soon observed that the Clarke generalized Jacobian can be replaced by another mapping whose elements fulfill a certain semismoothness property closely related to the one imposed on pseudosubgradients. This is particularly important for carrying over the semismooth Newton method to the infinite-dimensional setting, where no generalized Jacobian exists. We refer the reader to [31] for the semismooth Newton method in the infinite-dimensional setting. Though the semismoothness property imposed on the pseudosubgradients and the matrices used in the semismooth Newton method is essentially the same, there is a big difference: Whereas for the semismooth Newton method the semismoothness needs to hold only at the solution, for the pseudosubdiiferentials it has to hold at every $x\in\mathbb{R}^{n}$, which has far-reaching consequences.
Finally, in [7] the semismooth Newton method for solving inclusions
$0\in F(x)$
was introduced, where $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is a set-valued mapping. Here, the semismooth property was defined in terms of the limiting (Mordukhovich) coderivative and was later simplified to SC (subspace containing) derivative in [8]. The SC derivative can be considered as a generalization of the B-Jacobian (Bouligand) of single-valued mappings to multifunctions and acts as a kind of skeleton for the limiting coderivative. The semismooth Newton method has been already successfully applied to several difficult problems, see, e.g., [10, 11, 12].
We have seen that semismoothness plays an important role in solving different types of nonsmooth problems: optimization problems, equations and inclusions. The used notions of semismoothness differ slightly for the different problem types and so one aim of this paper is to introduce a unified notion which is compatible across the different problem types. This is important for the numerical solution of problems which are assembled by different nonsmooth objects. As an example consider an optimization problem of the form
$\min\,$ $\varphi(x,y)$
subject to $0\in F(x,y),$ (2)
$x\in\omega.$
where $\varphi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is a locally Lipschitz function, $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ is a set-valued mapping and $\omega\subset\mathbb{R}^{n}$ is a closed set. Problems of such type arise, e.g., in bilevel games and mathematical programs with equilibrium constraints (MPECs).
One possibility for tackling (2) is the so-called implicit programming approach (ImP), cf. [18]. Here we assume that for every $x$ the inclusion $0\in F(x,\cdot)$ has a unique solution denoted by $\sigma(x)$. Hence we can eliminate $y$ from the model above and arrive at the problem
$\min\,$ $\theta(x):=\varphi(x,\sigma(x))$
subject to $x\in\omega.$
Under the additional assumption that $\sigma$ is locally Lipschitz, the objective $\theta$ is locally Lipschitz as well and we can apply a bundle method for its numerical solution, provided a suitable oracle is available. We cannot guarantee that we actually compute a Clarke subgradient of $\theta$; we will show, however, that we can determine a pseudosubgradient, provided the original data $\varphi$ and $F$ are semismooth and semismooth, respectively.
In this paper we will provide the theoretical base for this approach: In particular, we will give all the necessary definitions and provide the basic calculus rules. We prove the semismooth property of the solution maps of inclusions as appearing in (2) by the semismooth property of the underlying mapping $F$. Up to now, only results on the semismooth property of single-valued solution mappings implicitely given by single-valued equations are available, cf. [30, 24, 13, 19, 16]
Moreover, we will also shed light on a remarkable property of graphically Lipschitzian semismooth set-valued mappings, namely that they are strictly proto-differentiable at almost all points of its graph. This is in accordance with the fact that semismooth single-valued maps are almost everywhere strictly differentiable.
Let us mention that for single-valued mappings there is a close relation between the semismoothness property and conservative derivatives, the latter being explored in a recent paper by Bolte and Pauwels [1]. In fact, it was shown in [6] that both notions are equivalent for semialgebraic maps. As semismooth derivatives also obey a chain rule like conservative Jacobians, the considerations about automatic differentiation from [1] carry over to our setting. Recently, a conservative Jacobian for solution maps given implicitely by equations was established in [2]. In [3], a conservative Jacobian for the solution map of a parametric monotone inclusion was obtained by rewriting the inclusion as an equivalent fixed-point equation for the forward-backward map by means of the resolvent. Though the resolvents are explicitly known for a lot of monotone mappings, it is not clear how a conservative Jacobian of the resolvent can be computed in the general case. In fact, the resolvent itself may be considered as the solution map of a shifted inclusion. On the contrary, our approach is based on the differentiation of the underlying set-valued mapping by SC derivatives.
Whereas this paper is of theoretical nature, in a subsequent paper we will demonstrate how we can apply the theory to several instances of problem (2) and will also provide some numerical experiments.
The paper is organized in the following way. Section 2 is subdivided into several subsections, the first two containing the basic notions of variational analysis and generalized differentiation. In Subsection 2.3 we describe a general principle, analogous to the construction of Clarke generalized Jacobian, how to generate from an arbitrary multifunction another multifunction having some favorable properties. In the next subsection we recall the basic properties of SCD (subspace containing derivative) mappings and give in Theorem 2.18 a characterization of strict proto-differentiability. In Subsection 2.5 we recall the semismooth property of set-valued mappings and its extension to SCD mappings.
Section 3 is devoted to the semismoothness property of single-valued mappings with respect to certain multifunctions. This concept is commonly used in semismooth Newton methods for solving nonsmooth equations, in particular in infinite dimensions. However, in contrast to these applications we do not consider the case that semismoothness only holds at the solution, but on the whole domain. There is a close relation with the seminal work of Norkin [22, 23] on generalized differentiation of scalar functions. In the next section we present some relations between the semismooth and the semismooth property for a class of multifunctions, whose graphs coincide locally after a change of coordinates with the one of a Lipschitzian mapping. Section 5 concerns the semismooth property of solution maps to inclusions. We establish the semismooth property of general solution maps and the semismooth property for continuous localizations and selections, respectively. We further demonstrate that the analysis considerably simplifies for SCD mappings.
The following notation is employed. Given a linear subspace $L\subseteq\mathbb{R}^{n}$, $L^{\perp}$ denotes its orthogonal complement. Further, given a multifunction $F$, $\mathrm{gph}\,F:=\{(x,y)\,\big{|}\,y\in F(x)\}$ stands for its graph. For an element $u\in\mathbb{R}^{n}$, $\|u\|$ denotes its Euclidean norm and ${\cal B}_{\delta}(u)$ denotes the closed ball around $u$ with radius $\delta$. The closed unit ball is denoted by ${\cal B}$. In a product space we use the norm $\|(u,v)\|:=\sqrt{\|u\|^{2}+\|v\|^{2}}$. The space of all $m\times n$ matrices is denoted by $\mathbb{R}^{m\times n}$. For a matrix $A\in\mathbb{R}^{m\times n}$ , we employ the operator norm $\|A\|$ with respect to the Euclidean norm and we denote the range of A by ${\rm rge\;}A$. Given a set $\Omega\subset\mathbb{R}^{s}$, we define the distance of a point $x$ to $\Omega$ by ${\rm dist}(x,\Omega):=\inf\{\|y-x\|\,\big{|}\,y\in\Omega\}$ and the indicator function is denoted by $\delta_{\Omega}$. Further we define for a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, and $x\in\Omega$ the moduli
${\rm lip\,}F(x):=\limsup_{{\textstyle{u,v\mathop{\rightarrow}\limits^{\Omega}x\atop u\not=v}}}\frac{\|F(u)-F(v)\|}{\|u-v\|},\quad\mathrm{clm}\,F(x):=\limsup_{u\mathop{\rightarrow}\limits^{\Omega}x}\frac{\|F(u)-F(x)\|}{\|u-x\|}.$
Clearly, $F$ is Lipschitz near $x$ (calm at $x$) if and only if ${\rm lip\,}F(x)$ ($\mathrm{clm}\,F(x)$) is finite.
2 Background from variational analysis and preliminary results
2.1 Nonsmooth analysis
Let us recall the notions of upper (outer) and lower (inner) limits in the sense of Painlevé–Kuratowski of a parameterized family $C_{t}$ of subsets of a metric space, where $t$ also belongs to a metric space:
$\mathop{{\rm Lim}\,{\rm sup}}_{t\to\bar{t}}C_{t}:=\{x\,\big{|}\,\liminf_{t\to\bar{t}}\,{\rm dist}(x,C_{t})=0\},$
$\mathop{{\rm Lim}\,{\rm inf}}_{t\to\bar{t}}C_{t}:=\{x\,\big{|}\,\limsup_{t\to\bar{t}}\,{\rm dist}(x,C_{t})=0\}.$
It easily follows from the definition that both the upper and lower limit sets are closed. These sets can also be described in terms of sequences as follows. The upper limit can be defined as the set of points $x$ for which there exists a sequence $t_{k}\to\bar{t}$ such that $x_{k}\to x$ for some $x_{k}\in C_{t_{k}}$ $\forall k$. Similarly, the lower limit can be defined as a set of points $x$ such that for every sequence $t_{k}\to\bar{t}$ we can find $x_{k}\in C_{t_{k}}$ such that $x_{k}\to x$. Clearly, $\mathop{{\rm Lim}\,{\rm inf}}_{t\to\bar{t}}C_{t}\subset\mathop{{\rm Lim}\,{\rm sup}}_{t\to\bar{t}}C_{t}$.
Consider a mapping $F:\Omega\to\mathbb{R}^{m}$, where $\Omega\subset\mathbb{R}^{n}$ is an open set. In the sequel we denote by ${\cal O}_{F}$ the set of points where $F$ is (Fréchet) differentiable and we denote the Jacobian by $\nabla F(x)$, $x\in{\cal O}_{F}$. We use this notation also for scalar-valued functions $f:\mathbb{R}^{n}\to\mathbb{R}$, i.e., $\nabla f(x)$ is an $1\times n$ matrix, i.e., a row vector.
The B-Jacobian (B-differential) is defined as
¯$\overline{\nabla}F(\bar{x})=\mathop{{\rm Lim}\,{\rm sup}}_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\{\nabla F(x)\},\ \bar{x}\in\Omega.$
Then the Clarke generalized Jacobian is given by $\mathrm{conv}\,\overline{\nabla}F(\bar{x})$.
If $F$ is locally Lipschitz, by Rademacher’s Theorem the set $\Omega\setminus{\cal O}_{F}$ is negligible, i.e., has Lebesgue measure 00. Since the Jacobians $\nabla F(x)$, $x\in{\cal O}_{F}$, are also locally bounded by the Lipschitz constant, we obtain ¯$\overline{\nabla}F(x)\subset\mathrm{conv}\,\overline{\nabla}F(x)$, $x\in\Omega$.
We say that $F$ is strictly differentiable at $\bar{x}\in\Omega$ if there is an $m\times n$ matrix $A$ such that
$\lim_{{\textstyle{x,x^{\prime}\to\bar{x}\atop x\not=x^{\prime}}}}\frac{F(x^{\prime})-F(x)-A(x^{\prime}-x)}{\|x^{\prime}-x\|}=0.$
Of course, this implies that $F$ is differentiable at $\bar{x}$ and $\nabla F(\bar{x})=A$. Strict differentiability of $F$ at $\bar{x}$ is equivalent with the requirements that $F$ is Lipschitz near $\bar{x}$ and
$\lim_{{\textstyle{x\to\bar{x}\atop t\downarrow 0}}}\frac{F(x+tu)-F(x)}{t}=Au,\ u\in\mathbb{R}^{n}$
for some $m\times n$ matrix $A$.
For a locally Lipschitzian function $f:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, we can define the Clarke subdifferential as
$\partial_{c}f(x):=\mathrm{conv}\,\overline{\nabla}f(x)^{T}:=\{g^{T}\,\big{|}\,g\in\mathrm{conv}\,\overline{\nabla}f(x)\},\ x\in\Omega.$
The elements of $\partial_{c}f(x)$ are called subgradients and are column vectors.
Given any extended real-valued function $f:\mathbb{R}^{n}\to(-\infty,\infty]$, we always assume that $f$ is proper, i.e., $f(x)>-\infty$ for all $x$ and $\mathrm{dom}\,f:=\{x\in\mathbb{R}^{n}\,\big{|}\,f(x)<\infty\}\not=\emptyset$. Given a point $\bar{x}\in\mathrm{dom}\,f$, the regular (Fréchet) subdifferential of $f$ at $\bar{x}$ is given by
^$\widehat{\partial}f(\bar{x}):=\Big{\{}x^{*}\in\mathbb{R}^{n}\,\big{|}\,\liminf_{x\to\bar{x}}\frac{f(x)-f(\bar{x})-\langle x^{*},x-\bar{x}\rangle}{\|x-\bar{x}\|}\geq 0\Big{\}},$
while the (limiting) subdifferential is defined by
$\partial f(\bar{x}):=\{x^{*}\,\big{|}\,\exists(x_{k},x_{k}^{*}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,\widehat{\partial}f}}(\bar{x},x^{*})\mbox{ with }f(x_{k})\to f(\bar{x})\}.$
If $\mathrm{dom}\,f=:\Omega$ is open, then clearly ${\cal O}_{f}\subset\mathrm{dom}\,\widehat{\partial}f$ and ¯$\overline{\nabla}f(\bar{x})^{T}\subset\partial f(\bar{x})$, provided that $f$ is continuous at $\bar{x}\in\Omega$. If $f$ is locally Lipschitz on $\Omega$ then
$\partial_{c}f(x)=\mathrm{conv}\,\partial f(x),\ x\in\Omega,$
cf. [28, Theorem 9.61].
Lemma 2.1.
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $F:\Omega\to\mathbb{R}^{m}$ be Lipschitz near $\bar{x}\in\Omega$. Then the following statements are equivalent:
  1. (i)
    $F$ is strictly differentiable at $\bar{x}$.
  2. (ii)
    ¯$\overline{\nabla}F(\bar{x})$ is a singleton.
  3. (iii)
    $\lim\limits_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists.
Proof.
ad (i) $\Rightarrow$ (iii): If $F$ is strictly differentiable at $\bar{x}$ then $\nabla F$ is continuous at $\bar{x}$ relative to ${\cal O}_{F}$, cf. [28, Exercise 9.64]. Thus $\lim\limits_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists (and equals to $\nabla F(\bar{x})$).
ad (iii) $\Rightarrow$ (ii): This is evident from the definitions.
ad (ii) $\Rightarrow$ (i): Let ¯$\overline{\nabla}F(\bar{x})=\{A\}$. For every $i=1,\ldots,m$, we conclude from [5, Theorem 2.6.6] that the $i$-th component of $F$, $i=1,\ldots m$, satisfies ¯$\bar{\nabla}F_{i}(\bar{x})=\{A_{i}\}$, where $A_{i}$ denotes the $i$-th row of $A$. Thus $\partial_{c}F_{i}(\bar{x})$ is a singleton and therefore $F_{i}$ is strictly differentiable at $\bar{x}$ by [5, Proposition 2.2.4]. Since every component $F_{i}$, $i=1,\ldots m$, is strictly differentiable at $\bar{x}$, so is $F$ as well. ∎
2.2 Variational geometry and differentiation of multifunctions
Definition 2.2.
Let $A$ be a set in $\mathbb{R}^{s}$ and let $\bar{x}\in A$. Then
  1. (i)
    The tangent (contingent, Bouligand) cone to $A$ at $\bar{x}$ is given by
    $T_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{t\downarrow 0}\frac{A-\bar{x}}{t}$
    and the paratingent cone to $A$ at $\bar{x}$ is given by
    $T^{P}_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{{\textstyle{x\mathop{\rightarrow}\limits^{A}\bar{x}\atop t\downarrow 0}}}\frac{A-x}{t}.$
    The regular (Clarke) tangent cone to $A$ at $\bar{x}$ amounts to
    $\widehat{T}_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{x\mathop{\rightarrow}\limits^{A}\bar{x}\atop t\downarrow 0}}}\frac{A-x}{t}.$
  2. (ii)
    The set
    $\widehat{N}_{A}(\bar{x}):=(T_{A}(\bar{x}))^{\circ},$
    i.e. the (negative) polar cone to the tangent cone $T_{A}(\bar{x})$, is the regular (Fréchet) normal cone to $A$ at $\bar{x}$, and
    $N_{A}(\bar{x}):=\mathop{{\rm Lim}\,{\rm sup}}\limits_{\stackrel{{\scriptstyle A}}{{x\rightarrow\bar{x}}}}\widehat{N}_{A}(x)$
    is the limiting (Mordukhovich) normal cone to $A$ at $\bar{x}$.
If $A$ is convex, then $\widehat{N}_{A}(\bar{x})=N_{A}(\bar{x})$ amounts to the classical normal cone in the sense of convex analysis and we will write $N_{A}(\bar{x})$.
The above listed cones enable us to describe the local behavior of set-valued maps via various generalized derivatives.
Definition 2.3.
Consider a multifunction $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ and let $(\bar{x},\bar{y})\in\mathrm{gph}\,F$.
  1. (i)
    The multifunction $DF(\bar{x},\bar{y}):\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ given by $\mathrm{gph}\,DF(\bar{x},\bar{y})=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is called the graphical derivative of $F$ at $(\bar{x},\bar{y})$.
  2. (ii)
    The multifunction $D_{*}F(\bar{x},\bar{y}):\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ given by $\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})=T^{P}_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is called the strict (paratingent) derivative of $F$ at $(\bar{x},\bar{y})$.
  3. (iii)
    The multifunction $\widehat{D}^{\ast}F(\bar{x},\bar{y}):\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{n}$ defined by
    $\mathrm{gph}\,\widehat{D}^{\ast}F(\bar{x},\bar{y})=\{(y^{*},x^{*})\,\big{|}\,(x^{*},-y^{*})\in\widehat{N}_{\mathrm{gph}\,F}(\bar{x},\bar{y})\}$
    is called the regular (Fréchet) coderivative of $F$ at $(\bar{x},\bar{y})$.
  4. (iv)
    The multifunction $D^{\ast}F(\bar{x},\bar{y}):\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{n}$, defined by
    $\mathrm{gph}\,D^{\ast}F(\bar{x},\bar{y})=\{(y^{*},x^{*})\,\big{|}\,(x^{*},-y^{*})\in N_{\mathrm{gph}\,F}(\bar{x},\bar{y})\}$
    is called the limiting (Mordukhovich) coderivative of $F$ at $(\bar{x},\bar{y})$.
Given a set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, we denote by $\mathrm{dom}\,F$ the set of all points $x$ with $F(x)\not=\emptyset$. We call $F$ single-valued, if for every $x\in\mathrm{dom}\,F$ the set $F(x)$ is a singleton. In this case, we can omit the second argument and write $DF(x)$, $\widehat{D}^{*}F(x),\ldots$ instead of $DF(x,F(x))$, $\widehat{D}^{*}F(x,F(x)),\ldots$. Given a mapping $F:\Omega\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$, we can identify it with a set-valued single-valued mapping $\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ via
$x\to\begin{cases}\{F(x)\}&\mbox{if $x\in\Omega$}\\ \emptyset&\mbox{else.}\end{cases}$
However, be aware that when considering limiting objects at $x$ where $F$ is not continuous, it is not enough to consider only sequences $x_{k}\to x$ but we must work with sequences $(x_{k},F(x_{k}))\to(x,F(x))$.
If a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is differentiable at $\bar{x}\in\Omega$, there holds
$DF(\bar{x})(u)=\nabla F(\bar{x})u,\ u\in\mathbb{R}^{n}.$
If $F$ is even strictly differentiable at $\bar{x}$, then also
$D_{*}F(\bar{x})(u)=\nabla F(\bar{x})u,\ u\in\mathbb{R}^{n},\quad D^{*}F(\bar{x})(v^{*})=\nabla F(\bar{x})^{T}v^{*},\ v^{*}\in\mathbb{R}^{m}$
holds. A generalization of strict differentiability to general set-valued mappings is provided by the notion of strict proto-differentiability.
Following [27], a set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is called proto-differentiable at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if the set-valued mappings
$\Delta_{\bar{x},\bar{y},t}F:\xi\mapsto\big{(}F(\bar{x}+t\xi)-\bar{y})/t\quad\mbox{ for $t>0$}$
graph-converge as $t\downarrow 0$, i.e., their graphs converge as subsets of $\mathbb{R}^{n}\times\mathbb{R}^{m}$. Of course, the limit must be equal to the graphical derivative $DF(\bar{x},\bar{y})$. This is the same as requiring that
$\mathop{{\rm Lim}\,{\rm sup}}_{t\downarrow 0}\frac{\mathrm{gph}\,F-(\bar{x},\bar{y})}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{t\downarrow 0}\frac{\mathrm{gph}\,F-(\bar{x},\bar{y})}{t}\big{(}=\mathrm{gph}\,DF(\bar{x},\bar{y})\big{)}.$
According to [25] we say that $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if, more generally, the mappings
$\Delta_{x,y,t}F:\xi\mapsto\big{(}F(x+t\xi)-y)/t\quad\mbox{ for $t>0$}$
graph-converge as $t\downarrow 0$ and $(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$, the limit being $D_{*}F(\bar{x},\bar{y})$. This means that,
$\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}\big{(}=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})\big{)}.$ (3)
Lemma 2.4.
If $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is strictly proto-differentiable at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ then $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and
$\mathrm{gph}\,DF(\bar{x},\bar{y})$ $=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})=\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\mathop{{\rm Lim}\,{\rm inf}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ (4)
$=\mathrm{gph}\,D_{*}F(\bar{x},\bar{y}).$
Proof.
By taking into account [28, Theorem 6.26], we obtain from (3) that
$\mathrm{gph}\,DF(\bar{x},\bar{y})$ $=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})\subset\mathrm{gph}\,D_{*}F(\bar{x},\bar{y})=\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}$
$=\mathop{{\rm Lim}\,{\rm inf}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\frac{\mathrm{gph}\,F-(x,y)}{t}=\mathop{{\rm Lim}\,{\rm inf}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})\subset T_{\mathrm{gph}\,F}(\bar{x},\bar{y}).$
Hence, in order to prove (4) there remains to show
$A:=\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)=\mathop{{\rm Lim}\,{\rm sup}}_{{\textstyle{t\downarrow 0\atop(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}}}\big{(}\mathrm{gph}\,F-(x,y)\big{)}/t=:B.$
Clearly, by the inclusions above we have $B=T_{\mathrm{gph}}\,F(\bar{x},\bar{y})\subset A$. Now consider $(u,v)\in A$ together with sequences $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$ and $(u_{k},v_{k})\in T_{\mathrm{gph}\,F}(x_{k},y_{k})$ satisfying $(u_{k},v_{k})\to(u,v)$. By definition of the tangent cone, for every $k$ we can find some $t_{k}\in(0,\frac{1}{k})$ and some $(\hat{x}_{k},\hat{y}_{k})\in\mathrm{gph}\,F$ with
$\|\frac{(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})}{t_{k}}-(u_{k},v_{k})\|\leq\frac{1}{k}$
implying $\big{(}(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\big{)}/t_{k}\to(u,v)$. This proves $(u,v)\in B$ and the equality $A=B$ is established.
Finally, in order to show that $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace, observe first that $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})=B=\widehat{T}_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a convex cone. Further, for any $(u,v)\in B=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ there are sequences $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$, $t_{k}\downarrow 0$ and $(\hat{x}_{k},\hat{y}_{k})\in T_{\mathrm{gph}\,F}(x_{k},y_{k})$ with $\big{(}(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\big{)}/t_{k}\to(u,v)$. Then there also holds $(\hat{x}_{k},\hat{y}_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})$ and
$-(u,v)=\lim_{k\to\infty}\frac{(x_{k},y_{k})-(\hat{x}_{k},\hat{y}_{k})}{t_{k}}\in B$
follows. Hence $B=-B$ and it follows that $B$ is a subspace. ∎
In case of single-valued Lipschitzian maps, strict proto-differentiability coincides with strict differentiability.
Lemma 2.5.
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $F:\Omega\to\mathbb{R}^{m}$ be Lipschitz near $\bar{x}\in\Omega$. Then $F$ is strictly proto-differentiable at $(\bar{x},F(\bar{x}))$ if and only if $F$ is strictly differentiable at $\bar{x}$.
Proof.
By Lipschitz continuity of $F$, for all convergent sequences $t_{k}\downarrow 0$, $u_{k}\to u$ and $x_{k}\to\bar{x}$ the sequence $(F(x_{k}+t_{k}u_{k})-F(x_{k}))/t_{k}$ has at least one accumulation point. Thus, strict proto-differentiability of $F$ at $\bar{x}$ is equivalent with the requirement that for such sequences $t_{k}$, $u_{k}$, $x_{k}$ the limits
$\lim_{k\to\infty}\frac{F(x_{k}+t_{k}u_{k})-F(x_{k})}{t_{k}}=\lim_{k\to\infty}\frac{F(x_{k}+t_{k}u)-F(x_{k})}{t_{k}}$
exist and equal to $D_{*}F(\bar{x})(u)$. Hence, if $F$ is strictly proto-differentiable at $(\bar{x},F(\bar{x}))$ then $D_{*}F(\bar{x})$ is single-valued, $\mathrm{dom}\,D_{*}F(\bar{x})=\mathbb{R}^{n}$ and $\mathrm{gph}\,D_{*}F(\bar{x})$ is a subspace by Lemma 2.4. We conclude that $D_{*}F(\bar{x},\bar{y})$ is in fact a linear mapping from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$ and therefore there exists an $m\times n$ matrix $A$ with $D_{*}F(\bar{x})(u)=Au$, $u\in\mathbb{R}^{n}$. Thus $F$ is strictly differentiable at $\bar{x}$. The reverse implication easily follows from the definition of strict differentiability and the fact that $D_{*}\nabla F(\bar{x})(u)=\nabla F(\bar{x})u$, $u\in\mathbb{R}^{n}$, in this case. ∎
2.3 Closure of mappings
It is well-known that the Clarke subdifferential $\partial_{c}\theta$ of a Lipschitz continuous function $\theta:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, has the following properties:
  • $\partial_{c}\theta$ is locally bounded
  • The set $\partial_{c}\theta(x)$ is nonempty, convex and compact for every $x\in\Omega$.
  • $\partial_{c}\theta$ is upper semicontinuous on $\Omega$.
We now discuss a possibility how to extend an arbitrary locally bounded mapping to a mapping having the properties stated above.
Given a mapping $\Psi:X\rightrightarrows Y$ between two metric spaces $X$ and $Y$, we denote by $\mathrm{cl}\,\Psi$ the mapping whose graph equals to the closure of $\mathrm{gph}\,\Psi$, i.e.,
$\mathrm{gph}\,(\mathrm{cl}\,\Psi)=\mathrm{cl}\,(\mathrm{gph}\,\Psi)=\{(x,y)\,\big{|}\,(x,y)=\lim_{k\to\infty}(x_{k},y_{k})\mbox{ for some sequence }(x_{k},y_{k})\in\mathrm{gph}\,\Psi\}.$
Obviously, $\mathrm{cl}\,(\mathrm{cl}\,\Psi)=\mathrm{cl}\,\Psi$ and for every $x\in\mathrm{dom}\,(\mathrm{cl}\,\Psi)$ we have that $(\mathrm{cl}\,\Psi)(x)$ is closed. Further, $\mathrm{dom}\,(\mathrm{cl}\,\Psi)\subset\mathrm{cl}\,(\mathrm{dom}\,\Psi)$ and equality holds, e.g., when the metric space $Y$ is compact.
In most cases we consider set-valued mappings between Euclidean spaces.
Definition 2.6.
Consider a mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$.
  1. 1.
    $\Psi$ is called locally bounded at $\bar{x}\in\mathbb{R}^{n}$ if there is a neighborhood $U$ of $\bar{x}$ and a nonnegative real $M$ such that
    $\|y\|\leq M\ \forall y\in\Psi(x)\ \forall x\in U.$ (5)
    We denote by $\mathrm{bnd}\,\Psi(\bar{x})$ the infimum of all reals $M$ such that (5) holds for some neighborhood $U$.
  2. 2.
    $\Psi$ is called outer semicontinuous (osc) at $\bar{x}\in\mathbb{R}^{n}$ if $(\mathrm{cl}\,\Psi)(\bar{x})=\Psi(\bar{x})$.
  3. 3.
    $\Psi$ is called upper semicontinuous (usc) at $\bar{x}$ if $\Psi(\bar{x})$ is closed and for every open set $V\supset\Psi(\bar{x})$ the set $\{x\,\big{|}\,\Psi(x)\subset V\}$ is a neighborhood of $\bar{x}$.
These terms are invoked on $X$, a subset of $\mathbb{R}^{n}$, when the property holds for every $x\in X$. The set $X$ will not be mentioned if $X=\mathbb{R}^{n}$.
By the definition, the function $x\mapsto\mathrm{bnd}\,\Psi(x)$ is usc.
Given $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, we denote by $\rm cocl\,\Psi$ the mapping defined by
$(\rm cocl\,\Psi)(x):=\mathrm{conv}\,\big{(}(\mathrm{cl}\,\Psi)(x)\big{)},\ x\in\mathbb{R}^{n}.$
It is easy to see that $\mathrm{dom}\,(\rm cocl\,\Psi)=\mathrm{dom}\,(\mathrm{cl}\,\Psi)$.
Lemma 2.7.
Let $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ be a mapping. If $\Psi$ is locally bounded at $x\in\mathbb{R}^{n}$ then $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ are also locally bounded at $x$ and $(\rm cocl\,\Psi)(x)$ is a convex and compact set. Further, both $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ are both osc and usc at $x$.
Proof.
Let $\Psi$ be locally bounded at $x\in\mathbb{R}^{n}$ and consider an open neighborhood $U$ of $x$ such that $V:=\bigcup_{x^{\prime}\in U}\Psi(x^{\prime})$ is bounded. We claim that
$\bigcup_{x^{\prime}\in U}(\mathrm{cl}\,{\Psi})(x^{\prime})\subset\bigcup_{x^{\prime}\in U}(\rm cocl\,{\Psi})(x^{\prime})\subset\mathrm{conv}\,(\mathrm{cl}\,V).$
Indeed, consider a point $(x^{\prime},y^{\prime})\in\mathrm{gph}\,(\mathrm{cl}\,\Psi)\cap(U\times\mathbb{R}^{m})$ together with a sequence $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,\Psi}}(x^{\prime},y^{\prime})$. Then for all $k$ sufficiently large we have $x_{k}\in U$ and consequently $y_{k}\in V$. Thus $y^{\prime}\in\mathrm{cl}\,V$ and $\bigcup_{x^{\prime}\in U}(\rm cocl\,{\Psi})(x^{\prime})\subset\mathrm{conv}\,(\mathrm{cl}\,V)$ follows. This verifies the local boundedness of $\rm cocl\,\Psi$ at $x$. Since the set $(\mathrm{cl}\,\Psi)(x)$ is closed and bounded, it is compact and the same holds true for its convex hull $(\rm cocl\,\Psi)(x)$.
Due to the local boundedness of $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ at $x$, the mappings $\mathrm{cl}\,\Psi$ and $\rm cocl\,\Psi$ are osc if and only if they are usc, cf. [28, Theorem 5.19]. Clearly, $\mathrm{cl}\,\Psi$ is osc at $x$ and there remains to show that $\rm cocl\,\Psi$ is usc at $x$. If $x\not\in\mathrm{dom}\,(\rm cocl\,\Psi)=\mathrm{dom}\,(\mathrm{cl}\,\Psi)$ then the set $\{x^{\prime}\,\big{|}\,x^{\prime}\not\in\mathrm{dom}\,(\mathrm{cl}\,\Psi)=\mathrm{dom}\,(\rm cocl\,\Psi)\}$ is a neighborhood of $x$ since $\mathrm{cl}\,\Psi$ is usc at $x$. Now let $x\in\mathrm{dom}\,(\rm cocl\,\Psi)$ and consider an open set $V\supset(\rm cocl\,{\Psi})(x)$. Since $(\rm cocl\,{\Psi})(x)$ is compact, we can find some $\epsilon>0$ such that the convex open set $(\rm cocl\,{\Psi})(x)+\epsilon{\cal B}$ is contained in $V$. Since the mapping $\mathrm{cl}\,\Psi$ is usc, the set $U:=\{x^{\prime}\,\big{|}\,(\mathrm{cl}\,\Psi)(x^{\prime})\subset(\mathrm{cl}\,\Psi)(x)+\epsilon{\cal B}\}$ is a neighborhood of $x$. For every $x^{\prime}\in U$ we have
$(\rm cocl\,{\Psi})(x^{\prime})=\mathrm{conv}\,\big{(}(\mathrm{cl}\,\Psi)(x^{\prime})\big{)}\subset\mathrm{conv}\,\big{(}(\mathrm{cl}\,\Psi)(x)+\epsilon{\cal B}\big{)}=(\rm cocl\,{\Psi})(x)+\epsilon{\cal B}\subset V$
and the lemma is proved. ∎
Example 2.8.
Let $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, be locally Lipschitzian and consider the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ given by
$\Psi_{F}(x):=\begin{cases}\{\nabla F(x)\}&\mbox{if $x\in{\cal O}_{F}$}\\ \emptyset&\mbox{else.}\end{cases}$
Then for every $x\in\Omega$ there holds ¯$\overline{\nabla}F(x)=(\mathrm{cl}\,\Psi_{F})(x)$ and $\mathrm{conv}\,\overline{\nabla}F(x)=(\rm cocl\,\Psi_{F})(x)$.
2.4 SCD mappings
In this subsection we collect only some basic facts from the theory of SCD-mappings which are needed in this paper. We refer the interested reader to [8, 9] for more details and more far-reaching properties.
Let us denote by ${\cal Z}_{nm}$ the metric space of all n-dimensional subspaces of $\mathbb{R}^{n+m}$ equipped with the metric
$d_{{\cal Z}_{nm}}(L_{1},L_{2}):=\|P_{1}-P_{2}\|,$
where $P_{i}$ is the symmetric $(n+m)\times(n+m)$ matrix representing the orthogonal projection onto $L_{i},i=1,2$.
For notational reasons, we usually identify $\mathbb{R}^{n+m}$ with $\mathbb{R}^{n}\times\mathbb{R}^{m}$, i.e., for a subspace $L\in{\cal Z}_{nm}$ we denote its elements by $(u,v)$ instead of $\begin{pmatrix}u\\ v\end{pmatrix}$. Analogously, given two matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{m\times n}$ so that $L={\rm rge\;}\begin{pmatrix}A\\ B\end{pmatrix}$, we write $L={\rm rge\;}(A,B)$.
Lemma 2.9.
cf.[9, Lemma 3.1] The metric space ${\cal Z}_{nm}$ is (sequentially) compact.
With each $L\in{\cal Z}_{nm}$ one can associate its adjoint subspace $L^{*}$ defined by
$L^{*}:=\{(-v^{*},u^{*})\in\mathbb{R}^{m}\times\mathbb{R}^{n}\,\big{|}\,(u^{*},v^{*})\in L^{\perp}\}.$
Since $\dim L^{\perp}=m$, it follows that $L^{*}\in{\cal Z}_{mn}$ (i.e., its dimension is $m$). It is easy to see that
$L^{*}=S_{nm}L^{\perp},\mbox{ where }S_{nm}=\left(\begin{array}[]{lc}0&-I_{m}\\ I_{n}&0\end{array}\right).$
Further we have $(L^{*})^{*}=L$ and $d_{{\cal Z}_{mn}}(L_{1}^{*},L_{2}^{*})=d_{{\cal Z}_{nm}}(L_{1},L_{2})$ for all subspaces $L_{1},L_{2}\in{\cal Z}_{nm}$ and thus the mapping $L\to L^{*}$ is an isometry between ${\cal Z}_{nm}$ and ${\cal Z}_{mn}$, cf. [9]. In what follows, the symbol $L^{*}$ signifies both the adjoint subspace to some $L\in{\cal Z}_{nm}$ as well as an arbitrary subspace from ${\cal Z}_{mn}$. This double role, however, cannot lead to a confusion.
Consider now a mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$.
Definition 2.10.
We say that $F$ is graphically smooth of dimension $n$ at $(\bar{x},\bar{y})$ if $T_{\mathrm{gph}\,F}(\bar{x},\bar{y})\in{\cal Z}_{nm}$. By $\mathcal{O}_{F}$ we denote the subset of $\mathrm{gph}\,F$, where $F$ is graphically smooth of dimension $n$.
Clearly, for $(x,y)\in\mathcal{O}_{F}$ and $L=T_{\mathrm{gph}\,F}(x,y)=\mathrm{gph}\,DF(x,y)$ it holds that $L^{\perp}=\hat{N}_{\mathrm{gph}\,F}(x,y)$ and $L^{*}=\mathrm{gph}\,\widehat{D}^{*}F(x,y)$.
Next we introduce the two derivative-like mappings ${\mathcal{S}}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows{\cal Z}_{nm}$ and ${\mathcal{S}}^{*}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows{\cal Z}_{mn}$ defined by
${\mathcal{S}}F(x,y):$ =$=\{L\in{\cal Z}_{nm}\,\big{|}\,\exists(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(x,y)\mbox{ such that }\lim_{k\to\infty}d_{{\cal Z}_{nm}}\big{(}L,T_{\mathrm{gph}\,F}(x_{k},y_{k})\big{)}=0\},$
and
${\mathcal{S}^{*}}F(x,y):$ =$=\{L^{*}\,\big{|}\,L\in{\mathcal{S}}F(x,y)\}.$
Both ${\mathcal{S}}F$ and ${\mathcal{S}}^{*}F$ are in fact generalized derivatives of $F$ whose elements, by virtue of the above definitions, are subspaces. Further we have the relation
$L^{*}\subset\mathrm{gph}\,D^{*}F(x,y)\ \forall L^{*}\in{\mathcal{S}}^{*}F(x,y),$
cf. [9, Eq 15].
In what follows ${\mathcal{S}}F$ will be called primal SC (subspace containing) derivative or SC limiting graphical derivative and ${\mathcal{S}}^{*}F$ will be termed adjoint SC derivative (SC limiting coderivative). Let us mention that in [9] the SC adjoint derivative ${\mathcal{S}}^{*}F$ was defined in a different but equivalent manner.
Note that ${\mathcal{S}}F$ coincide with the mapping $\mathrm{cl}\,\widehat{\mathcal{S}}F$, where
$\widehat{\mathcal{S}}F(x,y):=\begin{cases}T_{\mathrm{gph}\,F}(x,y)&\mbox{if $(x,y)\in{\cal O}_{F}$},\\ \emptyset&\mbox{else.}\end{cases}$
In particular, $\mathrm{gph}\,{\mathcal{S}}F$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{m}\times{\cal Z}_{nm}$ and, due to the isometry $L\leftrightarrow L^{*}$, $\mathrm{gph}\,{\mathcal{S}}^{*}F$ is closed in $\mathbb{R}^{n}\times\mathbb{R}^{m}\times{\cal Z}_{mn}$ as well.
On the basis of these SC derivatives we may now introduce the following notion playing a crucial role in the sequel.
Definition 2.11.
A mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is said to have the SCD property at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$, provided ${\mathcal{S}}^{*}F(\bar{x},\bar{y})\neq\emptyset$. $F$ is termed an SCD mapping if it has the SCD property at all points of $\mathrm{gph}\,F$.
By virtue of the definition, the SCD property at $(\bar{x},\bar{y})$ is obviously equivalent with the condition ${\mathcal{S}}F(\bar{x},\bar{y})\not=\emptyset$.
Since we consider convergence in the compact metric space ${\cal Z}_{nm}$, by using similar arguments as in the proof of [8, Lemma 3.6], we readily obtain the following result.
Lemma 2.12.
A mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ has the SCD property at $(x,y)\in\mathrm{gph}\,F$ if and only if $(x,y)\in\mathrm{cl}\,{\cal O}_{F}$. Further, $F$ is an SCD mapping if and only if $\mathrm{cl}\,{\cal O}_{F}=\mathrm{cl}\,\mathrm{gph}\,F$, i.e., $F$ is graphically smooth of dimension $n$ at the points of a dense subset of its graph.
For SCD mappings, the adjoint SC derivative ${\mathcal{S}}^{*}F$ forms some kind of skeleton for the limiting coderivative $D^{*}F$. It contains the information which is important in applications, in particular for numerical computations, but is much easier to compute than the limiting coderivative. On the other hand, the limiting coderivative is a more general object important for arbitrary set-valued mappings.
The derivatives ${\mathcal{S}}F$ and ${\mathcal{S}}^{*}F$ can be considered as a generalization of the B-Jacobian to multifunctions. In case of single-valued continuous mappings one has the following relationship.
Lemma 2.13 (cf.[9, Lemma 3.5]).
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $f:\Omega\to\mathbb{R}^{m}$ be continuous. Then for every $x\in\Omega$ there holds
${\mathcal{S}}f(x):={\mathcal{S}}\big{(}x,f(x)\big{)}\supseteq\{{\rm rge\;}(I,A)\,\big{|}\,A\in\overline{\nabla}f(x)\},$
${\mathcal{S}}^{*}f(x):={\mathcal{S}}^{*}\big{(}x,f(x)\big{)}\supseteq\{{\rm rge\;}(I,A^{T})\,\big{|}\,A\in\overline{\nabla}f(x)\}.$
If $f$ is Lipschitz near $x$, these inclusions hold with equality and $f$ has the SCD property at all points $x^{\prime}$ sufficiently close to $x$.
For the sake of completeness we state also the following calculus rules.
Proposition 2.14 (cf.[9, Theorem 4.1]).
Given a mapping $G:\mathbb{R}^{l}\rightrightarrows\mathbb{R}^{m}$ and a mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{l}\times\mathbb{R}^{m}$, consider the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ given by
$\mathrm{gph}\,F=\{(x,y)\,\big{|}\,\Phi(x,y)\in\mathrm{gph}\,G\}.$
Further assume that we are given $(x,y)\in\mathrm{gph}\,F$ such that $\Phi$ is continuously differentiable in some neighborhood of $(x,y)$, $\nabla\Phi(x,y)$ has full row rank and $G$ has the SCD property at $\Phi(x,y)$. Then $F$ has the SCD property at $(x,y)$ and
${\mathcal{S}}F(x,y)=\{L\in{\cal Z}_{nm}\,\big{|}\,\nabla\Phi(x,y)L\in{\mathcal{S}}G(\Phi(x,y))$
${\mathcal{S}}^{*}F(x,y)=S_{nm}\nabla\Phi(x,y)^{T}S_{lm}^{T}{\mathcal{S}}^{*}G(\Phi(x,y)).$
Proposition 2.15.
Let $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ have the SCD property at $(x,y)\in\mathrm{gph}\,F$ and let $h:\Omega\to\mathbb{R}^{m}$ be continuously differentiable at $x\in\Omega$ where $\Omega\subseteq\mathbb{R}^{n}$ is open. Then $F+h$ has the SCD property at $(x,y+h(x))$ and
${\mathcal{S}}(F+h)(x,y+h(x))=\Big{\{}\left(\begin{matrix}I_{n}&0\\ \nabla h(x)&I_{m}\end{matrix}\right)L\,\big{|}\,L\in{\mathcal{S}}F(x,y)\Big{\}},$
${\mathcal{S}}^{*}(F+h)(x,y+h(x))=\Big{\{}\left(\begin{matrix}I_{m}&0\\ \nabla h(x)^{T}&I_{n}\end{matrix}\right)L^{*}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,y)\Big{\}}.$
Proof.
Follows from Proposition 2.14 and the observation that $\mathrm{gph}\,(F+h)=\{(x,y)\,\big{|}\,(x,y-h(x))\in\mathrm{gph}\,F\}$. ∎
An important class of SCD mappings is constituted by graphically Lipschitzian mappings:
Definition 2.16 (cf.[28, Definition 9.66]).
A mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $d$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if there is an open neighborhood $W$ of $(\bar{x},\bar{y})$ and a one-to-one mapping $\Phi$ from $W$ onto an open subset of $\mathbb{R}^{n+m}$ with $\Phi$ and $\Phi^{-1}$ continuously differentiable, such that $\Phi(\mathrm{gph}\,F\cap W)$ is the graph of a Lipschitz continuous mapping $f:U\to\mathbb{R}^{n+m-d}$, where $U$ is an open set in $\mathbb{R}^{d}$.
From Proposition 2.14 and Lemma 2.13 one can now easily derive the following result:
Lemma 2.17 (cf.[9, Corollary 4.3]).
Assume that $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. Then $F$ has the SCD property at $(\bar{x},\bar{y})$. Moreover, if $\Phi$ and $f$ are as in Definition 2.16, then
${\mathcal{S}}F(\bar{x},\bar{y})=\{\nabla\Phi(\bar{x},\bar{y})^{-1}{\rm rge\;}(I,A)\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\},$
${\mathcal{S}}^{*}F(\bar{x},\bar{y})=\{S_{nm}\nabla\Phi(\bar{x},\bar{y})^{T}S^{T}_{nm}{\rm rge\;}(I,A^{T})\,\big{|}\,A\in\overline{\nabla}f(\bar{u})\},$
where $\big{(}\bar{u},f(\bar{u})\big{)}=\Phi(\bar{x},\bar{y})$.
Prominent examples for graphically Lipschitzian mappings are the subdifferential mapping $\partial f$ for prox-regular and subdifferentially continuous functions $f:\mathbb{R}^{n}\to(-\infty,\infty]$, cf. [28, Proposition 13.46], or more generally, locally maximally hypomonotone mappings, cf. [8].
Theorem 2.18.
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ has the SCD property at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. If $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ then ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton. Further, if $F$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})$ then the reverse implication also holds true.
Proof.
Assume that $F$ is strictly proto-differentiable, then we know by Lemma 2.4 that $\bar{L}:=T_{\mathrm{gph}\,F}(\bar{x},\bar{y})$ is a subspace. Further, from (4) we conclude that
$\bar{L}\supset\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(\bar{x},\bar{y})}T_{\mathrm{gph}\,F}(x,y)\supset\bigcup_{L\in{\mathcal{S}}F(\bar{x},\bar{y})}L\not=\emptyset.$
Hence, $\bar{n}:=\dim\bar{L}\geq n$ and if $\bar{n}=n$ then ${\mathcal{S}}F(\bar{x},\bar{y})=\{\bar{L}\}$ follows. Now assume that $\bar{n}>n$ and choose an orthogonal basis $(u_{i},v_{i})$, $i=1,\ldots,\bar{n}$, of $\bar{L}$. Since $(\bar{x},\bar{y})\in(\mathrm{cl}\,{\cal O}_{F})\setminus{\cal O}_{F}$, there must be a sequence $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}(\bar{x},\bar{y})$ satisfying $(x_{k},y_{k})\not=(\bar{x},\bar{y})$ $\forall k$. Taking into account (4), for every $k$ sufficiently large and every $i=1,\ldots,\bar{n}$ we can find elements $(u_{i}^{k},v_{i}^{k})\in L_{k}:=T_{\mathrm{gph}\,F}(x_{k},y_{k})$ with $\|(u_{i}^{k},v_{i}^{k})-(u_{i},v_{i})\|<1/\sqrt{\bar{n}}$. We claim now that for those $k$ the vectors $(u_{i}^{k},v_{i}^{k})$, $i=1,\ldots,\bar{n}$, are linearly independent. Assume on the contrary, that for some large $k$ these vectors are linearly dependent and we can find reals $\lambda_{i}^{k}$, $i=1,\ldots,\bar{n}$ with
$\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(u_{i}^{k},v_{i}^{k})=0\mbox{ and }\sum_{i=1}^{\bar{n}}(\lambda_{i}^{k})^{2}=1.$
Then by orthogonality of $(u_{i},v_{i})$, $i=1,\ldots,\bar{n}$, together with the Cauchy-Schwarz inequality we obtain that
11 11 $=\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(u_{i},v_{i})\|\leq\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}(u_{i}^{k},v_{i}^{k})\|+\|\sum_{i=1}^{\bar{n}}\lambda_{i}^{k}\big{(}(u_{i},v_{i})-(u_{i}^{k},v_{i}^{k})\big{)}\|$
$\leq\Big{(}\sum_{i=1}^{\bar{n}}(\lambda_{i}^{k})^{2}\Big{)}^{1/2}\Big{(}\sum_{i=1}^{\bar{n}}\|(u_{i},v_{i})-(u_{i}^{k},v_{i}^{k})\|^{2}\Big{)}^{1/2}<1\cdot(\sum_{i=1}^{\bar{n}}\frac{1}{\bar{n}})^{1/2}=1,$
a contradiction. Thus $(u_{i}^{k},v_{i}^{k})$, $i=1,\ldots,\bar{n}$, are linearly independent and $\dim L_{k}\geq\bar{n}$ follows. However, by definition of the set ${\cal O}_{F}$ we have $\dim L_{k}=n$, a contradiction. Hence $\bar{n}=n$ and ${\mathcal{S}}F(\bar{x},\bar{y})=\{\bar{L}\}$ holds true.
To show the second assertion, assume that $F$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})$ and let $W$, $\Phi$, $U$ and $f$ be as in Definition 2.16. Note that for all sequences $t_{k}\downarrow 0$, $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$ and $(\hat{x}_{k},\hat{y}_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$ such that
$\lim_{k\to\infty}\frac{(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})}{t_{k}}=:(\xi,\eta)$
exists, we have that $\Phi(x_{k},y_{k})=:(u_{k},f(u_{k})){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,f}}(\bar{u},f(\bar{u}))=\Phi(\bar{x},\bar{y})$, $\Phi(\hat{x}_{k},\hat{y}_{k})=:(\hat{u}_{k},f(\hat{u}_{k})){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,f}}(\bar{u},f(\bar{u}))$ and
$\lim_{k\to\infty}\frac{(\hat{u}_{k},f(u_{k}))-(u_{k},f(u_{k}))}{t_{k}}$ $=\lim_{k\to\infty}\frac{\nabla\Phi(\bar{x},\bar{y})\big{(}(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\big{)}+o(\|(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})\|)}{t_{k}}$
$=\nabla\Phi(\bar{x},\bar{y})(\xi,\eta).$
Analogously, for all sequences $t_{k}\downarrow 0$, $u_{k}{\mathop{\longrightarrow}\limits^{U}}\bar{u}$, $\hat{u}_{k}{\mathop{\longrightarrow}\limits^{U}}\bar{u}$ such that
$\lim_{k\to\infty}\frac{(\hat{u}_{k},f(\hat{u}_{k}))-(u_{k},f(u_{k}))}{t_{k}}=:(\sigma,\tau)$
exists, we have that $\Phi^{-1}(u_{k},f(u_{k}))=:(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$, $\Phi^{-1}(\hat{u}_{k},f(\hat{u}_{k}))=:(\hat{x}_{k},\hat{y}_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F\cap W}}(\bar{x},\bar{y})$ and
$\lim_{k\to\infty}\frac{(\hat{x}_{k},\hat{y}_{k})-(x_{k},y_{k})}{t_{k}}=\nabla\Phi(\bar{x},\bar{y})^{-1}(\sigma,\tau).$
Thus, $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if and only if $f$ is strictly proto-differentiable at $(\bar{u},f(\bar{u}))$, which is the same as $f$ being strictly differentiable at $\bar{u}$. Now, by Lemma 2.17 the SC derivative ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton if and only if ¯$\overline{\nabla}f(\bar{u})$ is a singleton, which is equivalent to strict differentiability of $f$ at $\bar{u}$. Combining these statements it follows that $F$ is strictly proto-differentiable at $(\bar{x},\bar{y})$ if and only if ${\mathcal{S}}F(\bar{x},\bar{y})$ is a singleton. ∎
2.5 On semismooth mappings
The following definition goes back to the recent paper [7].
Definition 2.19.
  1. 1.
    A set $A\subseteq\mathbb{R}^{n}$ is called semismooth at a point $\bar{x}\in A$ if for every $\epsilon>0$ there is some $\delta>0$ such that
    $|\langle x^{*},x-\bar{x}\rangle|\leq\epsilon\|x-\bar{x}\|\|x^{*}\|\ \forall x\in A\cap{\cal B}_{\delta}(\bar{x})\ \forall x^{*}\in N_{A}(x).$
  2. 2.
    A set-valued mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is called semismooth at a point $(\bar{x},\bar{y})\in\mathrm{gph}\,F$, if $\mathrm{gph}\,F$ is semismooth at $(\bar{x},\bar{y})$, i.e., for every $\epsilon>0$ there is some $\delta>0$ such that
    $|\langle x^{*},x-\bar{x}\rangle-\langle y^{*},y-\bar{y}\rangle|$
    $\qquad\leq\epsilon\|(x,y)-(\bar{x},\bar{y})\|\|(y^{*},x^{*})\|\ \forall(x,y)\in\mathrm{gph}\,F\cap{\cal B}_{\delta}(\bar{x},\bar{y})\ \forall(y^{*},x^{*})\in\mathrm{gph}\,D^{*}F(x,y).$ (6)
Note that the semismooth property was defined in [7] in a different but equivalent way. The class of semismooth mappings is rather broad. We mention here two important classes of multifunctions having this property.
Proposition 2.20 (cf.[8, Proposition 2.10]).
  1. (i)
    Every mapping whose graph is the union of finitely many closed convex sets is semismooth at every point of its graph.
  2. (ii)
    Every mapping with closed subanalytic graph is semismooth at every point of its graph.
For single-valued functions there holds the following result.
Proposition 2.21 (cf. [7, Proposition 3.7]).
Assume that $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a single-valued mapping which is Lipschitzian near $\bar{x}$. Then the following two statements are equivalent.
  1. (i)
    $F$ is semismooth at $\bar{x}$.
  2. (ii)
    For every $\epsilon>0$ there is some $\delta>0$ such that
    $\|F(x)-F(\bar{x})-C(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|\ \forall x\in{\cal B}_{\delta}(\bar{x})\ \forall C\in\mathrm{conv}\,\overline{\nabla}F(x),$
    i.e.,
    $\sup_{C\in\mathrm{conv}\,\overline{\nabla}F(x)}\|F(x)-F(\bar{x})-C(x-\bar{x})\|=o(\|x-\bar{x}\|).$ (7)
Let us compare the semismooth property with the semismooth property as introduced in [26].
Definition 2.22.
Let $\Omega\subseteq\mathbb{R}^{n}$ be nonempty and open. A function $F:\Omega\to\mathbb{R}^{m}$ is semismooth at $\bar{x}\in\Omega$, if it is Lipschitzian near $\bar{x}$ and if
$\lim_{{\textstyle{A\in\mathrm{conv}\,\overline{\nabla}F(\bar{x}+tu^{\prime})\atop u^{\prime}\to u,\ t\downarrow 0}}}Au^{\prime}$
exists for all $u\in\mathbb{R}^{n}$. If F is semismooth at all $\bar{x}\in\Omega$, we call $F$ semismooth on $\Omega$.
Then the following relation is well-known, see, e.g., [31, Proposition 2.7].
Proposition 2.23.
Let $F:\Omega\to\mathbb{R}^{m}$ be defined on the open set $\Omega\subset\mathbb{R}^{n}$. Then for $\bar{x}\in\Omega$ the following two statements are equivalent.
  1. 1.
    $F$ is semismooth at $\bar{x}$,
  2. 2.
    $F$ is Lipschitz near $\bar{x}$, the directional derivative $F^{\prime}(\bar{x};\cdot)$ exists and (7) is fulfilled.
Hence, by virtue of Proposition 2.21, for locally Lipschitz functions the difference between the semismooth property and the semismooth property is just the directional differentiability. Further, we want to mention that under local Lipschitz continuity the semismooth property as defined by Gowda [13] coincides with our semismooth property. However, there are also single-valued functions which are semismooth but not locally Lipschitz, e.g., the function $x\mapsto\sqrt{|x|}$ at 00. This example demonstrates that even for single-valued functions the semismooth property differs from the semismooth property more than only by the lack of directional differentiability.
Sometimes it is useful to replace $\mathrm{gph}\,D^{*}F(x,y)$ in the relation (6) by another set.
Definition 2.24.
Let $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ and $\Gamma:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}\times\mathbb{R}^{n}$ be mappings and let $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. We say that $F$ is semismooth with respect to $\Gamma$ (or shortly $\Gamma$-ss) at $(\bar{x},\bar{y})$, if the following three conditions are fulfilled:
  1. 1.
    $\Gamma$ is cone-valued,
  2. 2.
    there is some neighborhood $U$ of $(\bar{x},\bar{y})$ such that $\mathrm{gph}\,F\cap U\subset\mathrm{dom}\,\Gamma$,
  3. 3.
    for every $\epsilon>0$ there is some $\delta>0$ such that
    $|\langle x^{*},x-\bar{x}\rangle-\langle y^{*},y-\bar{y}\rangle|$
    $\qquad\leq\epsilon\|(x,y)-(\bar{x},\bar{y})\|\|(y^{*},x^{*})\|\ \forall(x,y)\in\mathrm{gph}\,F\cap{\cal B}_{\delta}(\bar{x},\bar{y})\ \forall(y^{*},x^{*})\in\Gamma(x,y).$
We say that $F$ is $\Gamma$-ss around $(\bar{x},\bar{y})$ if $F$ is $\Gamma$-ss for all $(x,y)\in\mathrm{gph}\,F$ sufficiently close to $(\bar{x},\bar{y})$. Finally, $F$ is called $\Gamma$-ss if $F$ is $\Gamma$-ss at every $(x,y)\in\mathrm{gph}\,F$.
Note that the semismooth property introduced in Definition Definition 2.19 amounts to the semismooth property of $F$ at $(\bar{x},\bar{y})$ with respect to $\Gamma(x,y)=\mathrm{gph}\,D^{*}F(x,y)=\{(y^{*},x^{*})\,\big{|}\,(x^{*},-y^{*})\in N_{\mathrm{gph}\,F}(x,y)\}$.
The notion of $\Gamma$-semismoothness has two main applications. Firstly, the calculus rules for coderivatives often yield only inclusions of the form $D^{*}F(x,y)(y^{*})\subset\Sigma(x,y)(y^{*})$. However, as in our motivating application, sometimes it is not necessary to compute the exact limiting coderivative but elements from $\Sigma$ related to the semismooth property. Thus, if we can show that $F$ is $\Gamma$-ss with respect to the mapping $\Gamma(x,y)=\mathrm{gph}\,\Sigma(x,y)$ then we can also work with $\Sigma$ instead of $D^{*}F$.
Secondly, the $\Gamma$-ss property will be useful in connection with the theory of SCD mappings.
Definition 2.25.
We say that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is SCD-ss at (around) $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ if it is semismooth at (around) $(\bar{x},\bar{y})$ with respect to the mapping $\bigcup{\mathcal{S}}^{*}F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}\times\mathbb{R}^{n}$ given by
$(\bigcup{\mathcal{S}}^{*}F)(x,y):=\bigcup_{L^{*}\in{\mathcal{S}}^{*}F(x,y)}L^{*}.$
The mapping $F$ is called SCD-ss if it is SCD-ss at every point of its graph.
It follows from the definition of the $\Gamma$-ss property that $F$ has the SCD property around $(\bar{x},\bar{y})$ whenever $F$ is SCD-ss at $(\bar{x},\bar{y})$.
The SCD-ss property has also a useful primal interpretation:
Proposition 2.26 (cf. [12, Proposition 2.12]).
Consider the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ and the pair $(\bar{x},\bar{y})\in\mathrm{gph}\,F$. Then the following two statements are equivalent:
  1. 1.
    $F$ is SCD-ss at $(\bar{x},\bar{y})$.
  2. 2.
    For every $\epsilon>0$ there is some $\delta>0$ such that for every $(x,y)\in\mathrm{gph}\,F\cap{\cal B}_{\delta}(\bar{x},\bar{y})$ and every $L\in{\mathcal{S}}F(x,y)$ there holds
    ${\rm dist}((x-\bar{x},y-\bar{y}),L)\leq\epsilon\|(x-\bar{x},y-\bar{y})\|.$
3 On $\Psi$-semismooth single-valued mappings
The following definition is closely related to the definition of semismooth operators in Banach spaces, cf. [31, Definition 3.1].
Definition 3.1.
Consider the mapping $F:\Omega\to\mathbb{R}^{m}$, where $\Omega\subset\mathbb{R}^{n}$ is open, and the multifunction $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$.
  1. 1.
    We say that $F$ is semismooth with respect to $\Psi$ (shortly, $\Psi$-ss) at $\bar{x}\in\Omega$ if $\mathrm{dom}\,\Psi$ is a neighborhood of $\bar{x}$, $\Psi$ is locally bounded at $\bar{x}$ and
    $\lim_{{\textstyle{x\mathop{\rightarrow}\limits^{\Omega}\bar{x}\atop x\not=\bar{x}}}}\frac{\sup\{\|F(x)-F(\bar{x})-A(x-\bar{x})\|\,\big{|}\,A\in\Psi(x)\}}{\|x-\bar{x}\|}=0.$
  2. 2.
    We say that $F$ is $\Psi$-ss on a subset $U\subset\Omega$ if $F$ is $\Psi$-ss at every $x\in U$.
Remark 3.2.
If $F$ is $\Psi$-ss at $\bar{x}$ then $\Psi$ is also called a Newton map for $F$ at $\bar{x}$ in [15]. If $\Psi$ is single-valued then it is also called a Newton derivative (or slanting function) for $F$ at $\bar{x}$ in the literature. We dispense with using this terminology because they are normally used at single points, namely solutions of the nonsmooth equation $F(x)=0$. However, for our purposes it is essential that the properties holds not only at one single point $\bar{x}$ but also for all $x$ belonging to some neighborhood of $\bar{x}$. In the literature one can find also several papers about the $\Psi$-ss property under different names using some additional assumptions on $\Psi$, see, e.g., [13], [22], [23].
Clearly, continuously differentiable mappings $F$ are $\nabla F$-ss.
Lemma 3.3.
Assume that $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is semismooth at $\bar{x}\in\Omega$ with respect to some mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$. Then $F$ is calm at $\bar{x}$ and $\mathrm{clm}\,F(\bar{x})\leq\mathrm{bnd}\,F(\bar{x})$.
Proof.
By definition, for every $\epsilon>0$ we can find some neighborhood $U\subset\mathrm{dom}\,\Psi(\bar{x})\cap\Omega$ of $\bar{x}$ such that for every $x\in U$ and every $A\in\Psi(x)$ there holds $\|A\|\leq\mathrm{bnd}\,\Psi(\bar{x})+\epsilon$ and $\|F(x)-F(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|$ implying $\|F(x)-F(\bar{x})\|\leq(\|A\|+2\epsilon)\|x-\bar{x}\|$. Since $\Psi(x)\not=\emptyset$ and $\epsilon>0$ can be made arbitrarily small, the assertion follows. ∎
It is well-known that Newton maps satisfy a chain rule, see, e.g., [15, Theorem 6.14]. For the sake of completeness we state the following result, where in contrast to [15] we do not assume Lipschitz continuity of the involved mappings $F_{1},F_{2}$:
Lemma 3.4.
Consider the mappings $F_{1}:U\to\mathbb{R}^{p}$, $F_{2}:V\to\mathbb{R}^{m}$, where $U\subset\mathbb{R}^{n}$ and $V\subset\mathbb{R}^{p}$ are open sets with $F_{1}(U)\subset V$. Assume that $F_{1}$ is semismooth at $\bar{x}\in U$ with respect to the mapping $\Psi_{1}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{p\times n}$ and that $F_{2}$ is semismooth at $F_{1}(\bar{x})$ with respect to the mapping $\Psi_{2}:\mathbb{R}^{p}\rightrightarrows\mathbb{R}^{m\times p}$. Then the composite mapping $F:U\to\mathbb{R}^{m}$ given by $F(x):=F_{2}(F_{1}(x))$ is $\Psi$-ss at $\bar{x}$ with $\Psi$ given by
$\Psi(x):=\{BA\,\big{|}\,A\in\Psi_{1}(x),\ B\in\Psi_{2}(F_{1}(x))\}.$
Proof.
By Lemma 3.3 we can find a neighborhood $\tilde{V}\subset V$ of $F_{1}(\bar{x})$ together with some real $\kappa_{2}>0$ such that $\|F_{2}(y)-F_{2}(F_{1}(\bar{x}))\|\leq\kappa_{2}\|y-F_{1}(\bar{x})\|$ and $\mathrm{bnd}\,\Psi_{2}(y)\leq\kappa_{2}$ $\forall y\in\tilde{V}$. Again by Lemma 3.3, we can find some neighborhood $\tilde{U}\subset U$ of $\bar{x}$ with $F_{1}(\tilde{U})\subset\tilde{V}$ together with $\kappa_{1}>0$ such that $\|F_{1}(x)-F_{1}(\bar{x})\|\leq\kappa_{1}\|x-\bar{x}\|$ and $\mathrm{bnd}\,\Psi_{1}(x)\leq\kappa_{1}$ for all $x\in\tilde{U}$. Next, for arbitrary $\epsilon>0$ we choose a neighborhood $V_{\epsilon}\subset\tilde{V}$ of $F_{1}(\bar{x})$ such that for each $y\in V_{\epsilon}$ and each $B\in\Psi_{2}(y)$ one has $\|F_{2}(y)-F_{2}(F_{1}(\bar{x}))-B(y-F_{1}(\bar{x}))\|\leq\epsilon\|y-F_{1}(\bar{x})\|$. Finally we choose a neighborhood $U_{\epsilon}\subset\tilde{U}$ such that $F_{1}(U_{\epsilon})\subset V_{\epsilon}$ and for every $x\in U_{\epsilon}$ and every $A\in\Psi_{1}(x)$ one has $\|F_{1}(x)-F_{1}(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|$. Thus, for every $x\in U_{\epsilon}$, every $A\in\Psi_{1}(x)$ and every $B\in\Psi_{2}(F_{1}(x))$ we obtain
$\|F(x)-F(\bar{x})-BA(x-\bar{x})\|$
$\leq\|F_{2}(F_{1}(x))-F_{2}(F_{1}(\bar{x}))-B(F_{1}(x)-F_{1}(\bar{x}))\|+\|B\big{(}F_{1}(x)-F_{1}(\bar{x})-A(x-\bar{x})\big{)}\|$
$\leq\epsilon\|F_{1}(x)-F_{1}(\bar{x})\|+\kappa_{2}\epsilon\|x-\bar{x}\|\leq\epsilon(\kappa_{1}+\kappa_{2})\|x-\bar{x}\|.$
In order to show the local boundedness of $\Psi$ at $\bar{x}$ just note that for every $x\in\tilde{U}$ and every $BA\in\Psi(x)$ we have $\|BA\|\leq\|B\|\|A\|\leq\kappa_{2}\kappa_{1}$. This proves that $F$ is $\Psi$-ss at $\bar{x}$. ∎
Using similar arguments one can prove the following lemma.
Lemma 3.5.
Consider an open set $\Omega\subset\mathbb{R}^{n}$, two mappings $F_{i}:\Omega\to\mathbb{R}^{m}$ and two mappings $\Psi_{i}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, $i=1,2$. If $F_{i}$ is $\Psi_{i}$-ss, $i=1,2$ at $\bar{x}\in\Omega$ then $F_{1}+F_{2}$ is $(\Psi_{1}+\Psi_{2})$-ss at $\bar{x}$.
The next auxiliary statement is important in the application we have in mind.
Lemma 3.6.
Let $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, be continuous and assume that $F$ is semismooth at $\bar{x}\in\Omega$ with respect to $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$. Then $F$ is also $\rm cocl\,\Psi$-ss at $\bar{x}$.
Proof.
Let $\epsilon>0$ be arbitrarily fixed. Since $\Psi$ is locally bounded at $\bar{x}$, we can find some $\bar{\delta}>0$ and some $M>0$ such that $\sup\{\|A\|\,\big{|}\,A\in\bigcup_{x\in{\cal B}_{\bar{\delta}}(\bar{x})}\Psi(x)\}\leq M$. By definition we can find some $\delta\in(0,\bar{\delta})$ such that for every $x\in{\cal B}_{\delta}(\bar{x})$ and every $A\in\Psi(x)$ there holds
$\|F(x)-F(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|.$
Now consider $x\in{\cal B}_{\delta/2}(\bar{x})\setminus\{\bar{x}\}$ and $A\in(\rm cocl\,\Psi)(x)$. Then $A$ can be written as a convex combination of elements in $\mathrm{cl}\,\Psi(x)$:
$A=\sum_{i=1}^{N}\lambda_{i}A_{i},\ \sum_{i=1}^{N}\lambda_{i}=1,\ \lambda_{i}\geq 0,\ A_{i}\in\mathrm{cl}\,\Psi(x),\ i=1,\ldots,N.$
For every matrix $A_{i}$ we can find some pair $(x_{i},B_{i})\in\mathrm{gph}\,\Psi$ with $\|A_{i}-B_{i}\|\leq\epsilon$, $\|x_{i}-x\|<\min\{\epsilon\|x-\bar{x}\|,\frac{\delta}{2}\}$ and $\|F(x)-F(x_{i})\|\leq\epsilon\|x-\bar{x}\|$. Then $x_{i}\in{\cal B}_{\delta}(\bar{x})$, $\|\bar{x}-x_{i}\|\leq(1+\epsilon)\|\bar{x}-x\|$ and
$\|F(x)-F(\bar{x})-A(x-\bar{x})\|=\|\sum_{i=1}^{N}\lambda_{i}\big{(}F(x)-F(\bar{x})-A_{i}(x-\bar{x})\big{)}\|\leq\sum_{i=1}^{N}\lambda_{i}\|F(x)-F(\bar{x})-A_{i}(x-\bar{x})\|$
$\leq\sum_{i=1}^{N}\lambda_{i}\big{(}\|F(x_{i})-F(\bar{x})-B_{i}(x_{i}-\bar{x})\|+\|F(x)-F(x_{i})\|+\|B_{i}(x_{i}-\bar{x})-A_{i}(x-\bar{x})\|\big{)}$
$\leq\sum_{i=1}^{N}\lambda_{i}\big{(}\epsilon\|x_{i}-\bar{x}\|+\epsilon\|x-\bar{x}\|+\|B_{i}\|\|x_{i}-x\|+\|B_{i}-A_{i}\|\|x-\bar{x}\|\big{)}$
$\leq\sum_{i=1}^{N}\lambda_{i}\big{(}\epsilon(1+\epsilon)\|x-\bar{x}\|+\epsilon\|x-\bar{x}\|+M\epsilon\|x-\bar{x}\|+\epsilon\|x-\bar{x}\|\big{)}=\epsilon(3+\epsilon+M)\|x-\bar{x}\|.$
Since $\epsilon>0$ can be taken arbitrary small, this inequality verifies that $F$ is $\rm cocl\,\Psi$-semismooth at $\bar{x}$. ∎
Consider now the following definition.
Definition 3.7 (cf.[22, 23]).
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is called generalized differentiable if there exists an usc mapping $G:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ with nonempty, convex and compact values $G(x)$ such that for every $\bar{x}\in\mathbb{R}^{n}$ there holds
$\sup_{g\in G(x)}|f(x)-f(\bar{x})-\langle g,x-\bar{x}\rangle|=o(\|x-\bar{x}\|).$
The elements of $G(x)$ are called pseudosubgradients.
By Lemma 3.6, a $\Psi$-ss function $f:\mathbb{R}^{n}\to\mathbb{R}$ is also generalized differentiable with
$G(x)=\{g\in\mathbb{R}^{n}\,\big{|}\,g^{T}\in(\rm cocl\,\Psi)(x)\}.$
Note that pseudosubgradients are column vectors whereas the elements of $\Psi$ are $1\times n$ matrices, i.e., row vectors.
Generalized differentiable functions have a lot of interesting properties, cf.[22, 23, 20], which can be carried over to $\Psi$-semismooth functions $f:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open. For the sake of completeness we state some of these properties in the following proposition.
Proposition 3.8.
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $f:\Omega\to\mathbb{R}$ be semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{1\times n}$. Then the following statements hold true.
  1. (i)
    For every $x\in\Omega$ there holds ^$\widehat{\partial}f(x)^{T}\subset(\rm cocl\,\Psi)(x)$.
  2. (ii)
    Let $x,x^{\prime}\in\Omega$ be such that the line segment $[x,x^{\prime}]$ connecting $x$ and $x^{\prime}$ is contained in $\Omega$. Then there exists $\bar{t}\in(0,1)$ and $g\in(\rm cocl\,\Psi)\big{(}x+\bar{t}(x^{\prime}-x)\big{)}$ with $f(x^{\prime})-f(x)=g(x^{\prime}-x)$.
  3. (iii)
    $f$ is locally Lipschitz on $\Omega$ and for $x\in\Omega$ there holds ${\rm lip\,}f(x)\leq\mathrm{bnd}\,\Psi(x)$.
  4. (iv)
    For all $x\in\Omega$ there holds $\partial f(x)^{T}\subset\partial_{c}f(x)^{T}\subset(\rm cocl\,\Psi)(x)$.
  5. (v)
    Let $\Omega_{\Psi}$ denote the set of all $x\in\Omega$ such that $(\rm cocl\,\Psi)(x)$ is a singleton. Then $\Omega\setminus\Omega_{\Psi}$ has Lebesgue measure zero. Moreover, the function $f$ is strictly differentiable at every $x\in\Omega_{\Psi}$ with $\{\nabla f(x)\}=(\rm cocl\,\Psi)(x)$ and the mapping $x\mapsto\nabla f(x)$ is continuous on $\Omega_{\Psi}$.
Proof.
ad (i): Assume on the contrary that there exist $x\in\Omega$ and $g^{T}\in\widehat{\partial}f(x)$ with $g\not\in(\rm cocl\,\Psi)(x)$ and consider the projection $\bar{g}$ of $g$ onto the convex compact set $(\rm cocl\,\Psi)(x)$, which must be different from $g$. Pick any sequence $t_{k}\downarrow 0$ and consider the points $x_{k}:=x+t_{k}(g-\bar{g})^{T}$. For sufficiently large $k$ we have $x_{k}\in\Omega$ and we can pick some $g_{k}\in\Psi(x_{k})$. Then the sequence $g_{k}$ is bounded and, by possibly passing to a subsequence, we can assume that $g_{k}$ converges to some $\hat{g}$. Clearly, $\hat{g}\in(\mathrm{cl}\,\Psi)(x)\subset(\rm cocl\,\Psi)(x)$ and therefore $(g-\bar{g})(\hat{g}-\bar{g})^{T}\leq 0$. Since $f$ is $\Psi$-ss on $\Omega$, we obtain that
$\lim_{k\to\infty}\frac{f(x_{k})-f(x)-g_{k}(x_{k}-x)}{\|x_{k}-x\|}=0.$
Further we have
$\liminf_{k\to\infty}\frac{f(x_{k})-f(x)-g(x_{k}-x)}{\|x_{k}-x\|}\geq 0$
due to $g^{T}\in\widehat{\partial}f(x)$ and consequently
$0\leq\liminf_{k\to\infty}\frac{(g_{k}-g)(x_{k}-x)}{\|x_{k}-x\|}=\frac{(\hat{g}-g)(g-\bar{g})^{T}}{\|g-\bar{g}\|}=\frac{(\hat{g}-\bar{g})(g-\bar{g})^{T}}{\|g-\bar{g}\|}-\|g-\bar{g}\|<0,$
a contradiction.
ad(ii): By Lemma 3.3, $f$ is calm at every $x\in\Omega$ and therefore continuous. Let :=${\cal I}:=\{t\in\mathbb{R}\,\big{|}\,x+t(x^{\prime}-x)\in\Omega\}\supset[0,1]$ and consider the function $\varphi:{\cal I}\to\mathbb{R}$, $\varphi(t):=f(x+t(x^{\prime}-x))-f(x)-t(f(x^{\prime})-f(x))$. Owing to Lemmata 3.4, 3.5, $\varphi$ is semismooth on ${\cal I}$ with respect to the mapping $\Phi(t):=\Psi(x+t(x^{\prime}-x))(x^{\prime}-x)-(f(x^{\prime})-f(x))$, $t\in{\cal I}$. There holds $\varphi(0)=\varphi(1)=0$ and, since $\varphi$ is continuous on $[0,1]$, there exists $\bar{t}\in(0,1)$ being either a global minimizer or a global maximizer of $\varphi$ on $[0,1]$. We may assume that $\bar{t}$ is a local minimizer, since otherwise we can interchange $x$ and $x^{\prime}$ to obtain a global minimizer $(1-\bar{t})$ for the function $-\varphi(1-t)$. Then
$0\in\widehat{\partial}\varphi(\bar{t})\subset(\rm cocl\,\Phi)(\bar{t})=(\rm cocl\,\Psi)(x+\bar{t}(x^{\prime}-x))(x^{\prime}-x)-(f(x^{\prime})-f(x))$
yielding the assertion.
ad(iii): This is an immediate consequence of (ii).
ad(iv): Since $f$ is locally Lipschitz on $\Omega$, for every $x\in\Omega$ there holds $\partial f(x)=(\mathrm{cl}\,\widehat{\partial}f)(x)$ by definition of $\partial f$ and $\partial_{c}f(x)=\mathrm{conv}\,\partial f(x)$ by [28, Theorem 9.61]. This yields, together with (i),
$\partial f(x)\subset\partial_{c}f(x)=(\rm cocl\,\widehat{\partial}f)(x)\subset\{g^{T}\,\big{|}\,g\in(\rm cocl\,(\rm cocl\,\Psi))(x)\}=\{g^{T}\,\big{|}\,g\in(\rm cocl\,\Psi)(x)\}.$
ad (v): In order to show that the set $\Omega\setminus\Omega_{\Psi}$ is negligible, one can take over the proof of [20, Theorem 1.12] to show that for every compact cube $K(x,\delta):=\{x^{\prime}\in\mathbb{R}^{n}\,\big{|}\,\max_{i}|x^{\prime}_{i}-x_{i}|\leq\delta\}\subset\Omega$ the set $K(x,\delta)\setminus\Omega_{\Psi}$ is negligible. Consider the countable collection $K(x_{1},\delta_{1}),K(x_{2},\delta_{2}),\ldots$ of cubes contained in $\Omega$ with rational $x_{i}$ and rational $\delta_{i}$. This collection covers $\Omega$ and therefore the set $\Omega\setminus\Omega_{\Psi}$ is negligible as the countable union of the negligible sets $K(x_{i},\delta_{i})\setminus\Omega_{\Psi}$.
For every $x\in\Omega_{\Psi}$ we conclude from (iv) that $\partial_{c}f(x)$ is a singleton and thus $f$ is strictly differentiable at $x$ with $\nabla f(x)=\{\partial_{c}f(x)\}$ by [5, Proposition 2.2.4]. Since strict differentiability of $f$ is equivalent with continuity of $\nabla f$ relative to ${\cal O}_{f}$, cf. [28, Exercise 9.64], $\nabla f$ is continuous on $\Omega_{\Psi}$. ∎
It is not difficult to extend these results to vector-valued mappings.
Proposition 3.9.
Let $\Omega\subset\mathbb{R}^{n}$ be open and assume that $F:\Omega\to\mathbb{R}^{m}$ is semismooth on $\Omega$ with respect to a mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$. Then the following statements hold:
  1. (i)
    $F$ is locally Lipschitz on $\Omega$.
  2. (ii)
    For every $x\in\Omega$ there holds $\mathrm{conv}\,\overline{\nabla}F(x)\subset(\rm cocl\,\Psi)(x)$.
  3. (iii)
    Let $\Omega_{\Psi}$ denote the set of all $x\in\Omega$ such that $(\rm cocl\,\Psi)(x)$ is a singleton. Then $\Omega\setminus\Omega_{\Psi}$ has Lebesgue measure zero. Moreover, the mapping $F$ is strictly differentiable at every $x\in\Omega_{\Psi}$ with $\{\nabla F(x)\}=(\rm cocl\,\Psi)(x)$ and the mapping $x\mapsto\nabla F(x)$ is continuous on $\Omega_{\Psi}$.
Proof.
ad (i): By Lemma 3.4 we obtain that each component function $F_{i}$, $i=1,\ldots,m$, is semismooth on $\Omega$ with respect to the mapping $\Psi_{i}:\Omega\rightrightarrows\mathbb{R}^{1\times n}$ given by $\Psi_{i}(x):=\{e_{i}^{T}A\,\big{|}\,A\in\Psi(x)\}$. Hence, the components $F_{i}$, $i=1,\ldots,m$ are locally Lipschitz on $\Omega$ and so is the mapping $F$ as well.
ad (ii): Let $\Pi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ denote the mapping which assigns to each $x$ the set of all $m\times n$ matrices whose $i$-th row belongs to $(\rm cocl\,\Psi_{i})(x)$. Clearly, $(\rm cocl\,\Psi)(x)\subset\Pi(x)$, $x\in\mathbb{R}^{n}$. Next let $\tilde{\Omega}:=\bigcap_{i=i}^{m}\Omega_{\Psi_{i}}$. For every $x\in\tilde{\Omega}$ we obtain from Proposition 3.8(v) that for every $i=1,\ldots,m$ the component function $F_{i}$ is strictly differentiable at $x$ and $(\rm cocl\,\Psi_{i})(x)=\{\nabla F_{i}(x)\}$ is a singleton. Hence, $F$ is also strictly differentiable at $x$ and $(\rm cocl\,\Psi)(x)=\Pi(x)=\{\nabla F(x)\}$. Since the set $\Omega\setminus\tilde{\Omega}=\bigcup_{i=1}^{m}(\Omega\setminus\Omega_{\Psi_{i}})$ is negligible, the set $\tilde{\Omega}$ differs from the set ${\cal O}_{F}$, the set of all points where $F$ is differentiable, by a negligible set. By denoting
$\nabla_{s}F(x):=\begin{cases}\nabla F(x)&\mbox{if $x\in\tilde{\Omega}$,}\\ \emptyset&\mbox{else}\end{cases}$
and taking into account $\mathrm{conv}\,\overline{\nabla}F=\rm cocl\,\nabla F$, it follows from [32, Theorem 4] that $\mathrm{conv}\,\overline{\nabla}F=\rm cocl\,\nabla_{s}F$. Since $\nabla_{s}F(x)\subset(\rm cocl\,\Psi)(x)$ $\forall x\in\Omega$, we obtain $\mathrm{conv}\,\overline{\nabla}F=(\rm cocl\,\nabla_{s}F)(x)\subset\big{(}\rm cocl\,(\rm cocl\,\Psi)\big{)}(x)=(\rm cocl\,\Psi)(x)$ $\forall x\in\Omega$.
ad (iii): Since $\tilde{\Omega}\subset\Omega_{\Psi}\subset\Omega$ and the set $\Omega\setminus\tilde{\Omega}$ is negligible, so must be the set $\Omega\setminus\Omega_{\Psi}$ as well. For every $x\in\Omega_{\Psi}$ we deduce from (ii) that $\mathrm{conv}\,\overline{\nabla}F(\bar{x})$ is also a singleton and therefore $F$ is strictly differentiable at $x$ by Lemma 2.1.
Finally, the continuity of $\nabla F$ on $\Omega_{\Psi}$ is again a consequence of [28, Exercise 9.64]. ∎
The following theorem is a reformulation of [4, Theorem 3.2] using our notation. Consider the problem
$\min f(x)\quad\mbox{subject to}\quad x\in\omega,$
where $f:\mathbb{R}^{n}\to\mathbb{R}$ and $\omega\subset\mathbb{R}^{n}$ is a convex polyhedron.
Theorem 3.10.
Assume that $f$ is semismooth with respect to some mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ and suppose that at each iteration of the BT-algorithm [33] instead of the oracle (1) one uses another
oracle, which returns for arbitrary $x\in\mathbb{R}^{n}$ the function value $f(x)$ and one element $g(x)$ with $g(x)^{T}\in\Psi(x)$.
Then, if the sequence $z_{k}$ produced by the BT-algorithm remains bounded, there exists an accumulation point $\bar{z}$ of $z_{k}$ such that $0\in(\rm cocl\,\Psi)(\bar{z})+N_{\omega}(\bar{z})$.
Note that the condition $0\in(\rm cocl\,\Psi)(\bar{z})+N_{\omega}(\bar{z})$ is in fact a necessary optimality condition for a local minimizer because by Proposition 3.8 we have $\partial_{c}f(\bar{z})\subset(\rm cocl\,\Psi)(\bar{z})$.
4 On the SCD-ss property of graphically Lipschitzian mappings and consequences
Proposition 4.1.
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and let $\Phi$ and $f$ be as in Definition 2.16. Then $F$ is SCD-ss at $(\bar{x},\bar{y})$ if and only if $f$ is ¯$\overline{\nabla}f$-semismooth at $\bar{u}$, where $(\bar{u},f(\bar{u}))=\Phi(\bar{x},\bar{y})$.
Proof.
Consider $(x,y)\in\mathrm{gph}\,F\cap W$ and the corresponding pair $(u,f(u))=\Phi(x,y)$, where $W$ is as in Definition 2.16. Let $L\in{\mathcal{S}}F(x,y)$ and $A\in\overline{\nabla}f(u)$ such that $L=\nabla\Phi(x,y)^{-1}({\rm rge\;}(I,A)$ according to Lemma 2.17. Then
${\rm dist}((x-\bar{x},y-\bar{y}),L)$ $=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}\big{(}\nabla\Phi(x,y)(x-\bar{x},y-\bar{y})-(p,Ap)\big{)}\|$
$=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}\big{(}\Phi(x,y)-\Phi(\bar{x},\bar{y})+o(\|(x-\bar{x},y-\bar{y})\|)-(p,Ap)\big{)}\|$
$\leq\|\nabla\Phi(x,y)^{-1}\|\big{(}\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|+o(\|(x-\bar{x},y-\bar{y})\|)\big{)}$
$\leq\|\nabla\Phi(x,y)^{-1}\|\big{(}\|f(u)-f(\bar{u})-A(u-\bar{u})\|+o(\|(x-\bar{x},y-\bar{y})\|)\big{)}.$
On the other hand there holds
$\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|^{2}=\inf_{q\in\mathbb{R}^{n}}\|q\|^{2}+\|f(u)-f(\bar{u})-A(u-\bar{u})-Aq)\|^{2}$
$\geq\inf_{w\in\mathbb{R}^{n}}\frac{\|w\|^{2}}{\|A\|^{2}}+\|f(u)-f(\bar{u})-A(u-\bar{u})-w)\|^{2}=\frac{1}{1+\|A\|^{2}}\|f(u)-f(\bar{u})-A(u-\bar{u})\|^{2},$
resulting in
${\rm dist}((x-\bar{x},y-\bar{y}),L)=\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}\big{(}\Phi(x,y)-\Phi(\bar{x},\bar{y})+o(\|(x-\bar{x},y-\bar{y})\|)-(p,Ap)\big{)}\|$
$\geq\inf_{p\in\mathbb{R}^{n}}\|\nabla\Phi(x,y)^{-1}(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|-\|\nabla\Phi(x,y)^{-1}\|o(\|(x-\bar{x},y-\bar{y})\|)$
$\geq\frac{1}{\|\nabla\Phi(x,y)\|}\inf_{p\in\mathbb{R}^{n}}\|(u-\bar{u}-p,f(u)-f(\bar{u})-Ap)\|-\|\nabla\Phi(x,y)^{-1}\|o(\|(x-\bar{x},y-\bar{y})\|)$
$\geq\frac{1}{\|\nabla\Phi(x,y)\|\sqrt{1+\|A\|^{2}}}\|f(u)-f(\bar{u})-A(u-\bar{u})\|-\|\nabla\Phi(x,y)^{-1}\|o(\|(x-\bar{x},y-\bar{y})\|).$
Since $\|A\|$ is bounded by the Lipschitz constant $\kappa$ of $f$ and
$\|u-\bar{u}\|\begin{cases}\geq\frac{\|(u-\bar{u},f(u)-f(\bar{u}))\|}{\sqrt{1+\kappa^{2}}}\geq\frac{\|(x-\bar{x},y-\bar{y})\|}{\|\nabla\Phi(x,y)^{-1}\|\sqrt{1+\kappa^{2}}}+o(\|(x-\bar{x},y-\bar{y})\|)&,\\ \leq\|(u-\bar{u},f(u)-f(\bar{u}))\|\leq\|\nabla\Phi(x,y)\|\|(x-\bar{x},y-\bar{y})\|+o(\|(x-\bar{x},y-\bar{y})\|),\end{cases}$
we conclude from these inequalities that
$\mathop{{\rm Lim}\,{\rm sup}}_{(x,y){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(\bar{x},\bar{y})}\sup_{L\in{\mathcal{S}}F(x,y)}\frac{{\rm dist}((x-\bar{x},y-\bar{y}),L)}{\|(x-\bar{x},y-\bar{y})\|}=0\ \Longleftrightarrow\ \limsup_{u\to\bar{u}}\sup_{A\in\overline{\nabla}f(u)}\frac{\|f(u)-f(\bar{u})-A(u-\bar{u})\|}{\|u-\bar{u}\|}=0$
and the assertion follows from Proposition 2.26 and the definition of ¯$\overline{\nabla}f$-semismoothness. ∎
Corollary 4.2.
Assume that the mapping $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ is graphically Lipschitzian of dimension $n$ at $(\bar{x},\bar{y})\in\mathrm{gph}\,F$ and is SCD-ss around $(\bar{x},\bar{y})$. Then there is an open neighborhood $W$ of $(\bar{x},\bar{y})$ such that for almost all $(x,y)\in\mathrm{gph}\,F\cap W$ (with respect to the $n$-dimensional Hausdorff measure) the mapping $F$ is strictly proto-differentiable at $(x,y)$.
Proof.
According to the notation of Definition 2.16, let $\Omega_{F}$ denote the set of all points in $\mathrm{gph}\,F\cap W$ where $F$ is strictly proto-differentiable and let $\Omega_{f}$ denote the set of all points $u\in U\subset\mathbb{R}^{n}$ where $f$ is strictly differentiable. We may also assume that $F$ is SCD-ss on $\mathrm{gph}\,F\cap W$ and therefore $f$ is ¯$\overline{\nabla}f$-ss on $U$ by Proposition 4.1. We conclude from Proposition 3.9 that the set $U\setminus\Omega_{f}$ has Lebesgue measure zero and the assertion follows from the relation
$\Omega_{F}=\Phi^{-1}(\{(u,f(u))\,\big{|}\,u\in\Omega_{f}\}),$
which holds true by Theorem 2.18 and Lemmata 2.1, 2.17. ∎
5 On the semismooth property of solution maps to inclusions
Theorem 5.1.
Consider a mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{l}$ and let
[…]
be the solution map of the inclusion $0\in F(x,y)$. Assume that we are given $(\bar{x},\bar{y})\in\mathrm{gph}\,S$ such that the qualification condition
$(0,0)\in D^{*}F(\bar{x},\bar{y},0)(z^{*})\ \Rightarrow\ z^{*}=0$ (8)
holds. Then there is a neighborhood $U$ of $(\bar{x},\bar{y})$ such that for all $(x,y)\in\mathrm{gph}\,S\cap U$ one has
$D^{*}S(x,y)(y^{*})\subset\Sigma(x,y)(y^{*}):=\{x^{*}\,\big{|}\,\exists z^{*}\in\mathbb{R}^{l}\mbox{ with }(x^{*},-y^{*})\in D^{*}F(x,y,0)(z^{*})\}.$ (9)
Moreover, $S$ is semismooth at $(x,y)$ with respect to the mapping $\Gamma:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}\times\mathbb{R}^{n}$ given by
$\Gamma(u,v)=\{(y^{*},x^{*})\,\big{|}\,x^{*}\in\Sigma(u,v)(y^{*})\},$
provided $F$ is semismooth at $(x,y,0)$.
Proof.
We claim that there is some open neighborhood $U$ of $(\bar{x},\bar{y})$ and some real $\gamma>0$ such that for every pair $(x,y)\in\mathrm{gph}\,S\cap U$, every $z^{*}\in\mathbb{R}^{l}$, $\|z^{*}\|=1$ and every $(x^{*},y^{*})\in D^{*}F(x,y,0)(z^{*})$ one has $\|(x^{*},y^{*})\|\geq 1/\gamma$. Assume on the contrary that there are sequences $(x_{k},y_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,S}}(\bar{x},\bar{y})$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in\mathrm{gph}\,D^{*}F(x_{k},y_{k},0)$ with $\|z_{k}^{*}\|=1$ and $\|(x_{k}^{*},y_{k}^{*})\|\to 0$ as $k$ tends to $\infty$. By possibly passing to a subsequence we can assume that $z_{k}^{*}$ converges to some $z^{*}$ with $\|z^{*}\|=1$ and by definition of the limiting coderivative we obtain $(0,0,z^{*})\in\mathrm{gph}\,D^{*}F(\bar{x},\bar{y},0)$ contradicting (8). Hence our claim holds true.
Then we also have for all $(x,y)\in\mathrm{gph}\,S\cap U$ that
$(0,0)\in D^{*}F(x,y,0)(z^{*})\ \Rightarrow\ z^{*}=0$
and the inclusion (9) follows from [21, Theorem 4.46].
Now assume that $F$ is semismooth at $(x,y,0)$ with $(x,y)\in U$. Fix $\epsilon>0$ and choose $\delta>0$ such that ${\cal B}_{\delta}(x,y)\subset U$ and for all $(x^{\prime},y^{\prime})\in\mathrm{gph}\,S\cap{\cal B}_{\delta}(x,y)$ and all $(z^{*},x^{*},y^{*})\in\mathrm{gph}\,D^{*}F(x^{\prime},y^{\prime},0)$ one has
$|\langle z^{*},0-0\rangle-\langle x^{*},x^{\prime}-x\rangle-\langle y^{*},y^{\prime}-y\rangle|\leq\epsilon\|(z^{*},x^{*},y^{*})\|\|(x^{\prime},y^{\prime},0)-(x,y,0)\|.$ (10)
Next consider $(y^{*},x^{*})\in\Gamma(x^{\prime},y^{\prime})=\mathrm{gph}\,\Sigma(x^{\prime},y^{\prime})$ and the corresponding $z^{*}$ verifying $(z^{*},x^{*},-y^{*})\in\mathrm{gph}\,D^{*}F(x^{\prime},y^{\prime},0)$. By our claim we have $\|z^{*}\|\leq\gamma\|(x^{*},-y^{*})\|=\gamma\|(x^{*},y^{*})\|$ and from (10) we obtain
$|\langle y^{*},y^{\prime}-y\rangle-\langle x^{*},x^{\prime}-x\rangle|\leq\epsilon\sqrt{1+\gamma^{2}}\|(x^{*},y^{*})\|\|(x^{\prime},y^{\prime})-(x,y)\|.$
Hence $S$ is $\Gamma$-ss at $(x,y)$ and the proof is complete. ∎
Remark 5.2.
This result extends slightly [21, Corollary 4.47], because estimate (9) holds true not only at the reference point $(\bar{x},\bar{y})$ but also on a neighborhood $U$ of it.
Theorem 5.3.
In the setting of Theorem 5.1, assume that there is an open neighborhood $\Omega\times V$ of $(\bar{x},\bar{y})\in\mathrm{gph}\,S$ such that $\mathrm{gph}\,S\cap(\Omega\times V)$ is the graph of a continuous mapping $\sigma:\Omega\to\mathbb{R}^{m}$. Further assume that the strengthened qualification condition
$(x^{*},0)\in D^{*}F(\bar{x},\bar{y},0)(z^{*})\ \Rightarrow z^{*}=0,x^{*}=0$ (11)
holds. Then there is some open neighborhood $U$ of $\bar{x}$ such that $\sigma$ is Lipschitz continuous on $U$. If, in addition, $F$ is semismooth at $(x,\sigma(x),0)$ for some $x\in U$, then for every function $\varphi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$, which is semismooth at $(x,\sigma(x))$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$, the mapping $\theta:U\to\mathbb{R}$, $\theta(u):=\varphi(u,\sigma(u))$, is semismooth at $x$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ given by
$\Psi(u):=\begin{cases}\{g_{x}^{T}+{x^{*}}^{T}\,\big{|}\,x^{*}\in\Sigma(u,\sigma(u))(g_{y}),(g_{x}^{T},g_{y}^{T})\in\Phi(u,\sigma(u))\}&\mbox{if $u\in U$,}\\ \emptyset&\mbox{else,}\end{cases}$
where $\Sigma$ is given by (9).
Proof.
We claim that there is some neighborhood $U\subset\Omega$ of $\bar{x}$ and some real $\gamma>0$ such that for all $x\in U$ and all $(z^{*},x^{*},y^{*})\in\mathrm{gph}\,D^{*}F(x,\sigma(x),0)$ one has
$\|y^{*}\|\geq\frac{1}{\gamma}\|(x^{*},z^{*})\|.$
Assume on the contrary that there are sequences $x_{k}\to\bar{x}$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in\mathrm{gph}\,D^{*}F(x_{k},\sigma(x_{k}),0)$ with $\|(x^{*}_{k},z^{*}_{k})\|=1$ and $y_{k}^{*}\to 0$ as $k\to\infty$. By possibly passing to a subsequence we can assume that $(x_{k}^{*},z_{k}^{*})$ converges to some $(x^{*},z^{*})$ with $\|(x^{*},z^{*})\|=1$. It follows that $(z^{*},x^{*},0)\in\mathrm{gph}\,D^{*}F(\bar{x},\sigma(\bar{x}),0)$ contradicting (11). Hence our claim holds true. Consequently, there holds
$(x^{*},0)\in D^{*}F(x,\sigma(x),0)(z^{*})\ \Rightarrow z^{*}=0,x^{*}=0$
for every $x\in U$. By possibly shrinking $U$ we may assume that $U$ is an open, bounded and convex neighborhood of $\bar{x}$. By invoking [21, Corollary 4.60] we obtain that $\sigma$ is Lipschitz near every $x\in U$ with
${\rm lip\,}\sigma(x)\leq\sup\{\|x^{*}\|\,\big{|}\,\exists z^{*}\mbox{ with }(z^{*},x^{*},-y^{*})\in\mathrm{gph}\,D^{*}F(x,\sigma(x),0),\|y^{*}\|\leq 1\}\leq\gamma,\ x\in U.$
Thus, by [28, Theorem 9.2], we obtain that $\sigma$ is Lipschitz continuous on $U$ with constant $\gamma$. Hence, by virtue of (9) we have
$D^{*}\sigma(x)(y^{*})\subset\Sigma(x,\sigma(x),0)\ \forall x\in U\forall y^{*}\in\mathbb{R}^{m}.$ (12)
Next consider $x\in U$ and assume that $F$ is semismooth at $(x,\sigma(x),0)$ and consider any function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ which is semismooth at $x$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$. Then $\Phi$ is locally bounded at $(x,\sigma(x))$ and we can find some open convex neighborhood $\tilde{U}\subset U$ of $x$ such that $\{(u,\sigma(u))\,\big{|}\,u\in\tilde{U}\}\subset\mathrm{dom}\,\Phi$ and
$M:=\sup\{\|(g_{x}^{T},g_{y}^{T})\|\,\big{|}\,(g_{x}^{T},g_{y}^{T})\in\Phi(u,\sigma(u)),u\in\tilde{U}\}<\infty.$
Further, for every $u\in U$ and every $y^{*}\in\mathbb{R}^{m}$ there holds
$\sup\{\|x^{*}\|\,\big{|}\,x^{*}\in\Sigma(u,\sigma(u))(y^{*})\}=\sup\{\|x^{*}\|\,\big{|}\,\exists z^{*}\mbox{ with }(z^{*},x^{*},-y^{*})\in D^{*}F(u,\sigma(u),0)\}\leq\gamma\|y^{*}\|.$
Hence, for every $u\in\tilde{U}$ and every ${u^{*}}^{T}\in\Psi(u)$ we have
$\|u^{*}\|\leq\sup\{\|g_{x}\|+\|x^{*}\|\,\big{|}\,x^{*}\in\Sigma(u,\sigma(u))(g_{y}),\ (g_{x}^{T},g_{y}^{T})\in\Phi(u,\sigma(u))\}\leq M+\gamma M,$
verifying the local boundedness of $\Psi$ at $x$. Further, since $\mathrm{dom}\,\Phi$ is a neighborhood of $(x,\sigma(x))$ and taking into account (12), we obtain that $\mathrm{dom}\,\Psi$ is a neighborhood of $x$. Finally, let $\epsilon>0$ be given and consider $\delta>0$ such that for all $x^{\prime}\in U$ with $(x^{\prime},\sigma(x^{\prime}),0)\in{\cal B}_{\delta}(x,\sigma(x),0)$ and all $(z^{*},x^{*},y^{*})\in\mathrm{gph}\,D^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ one has
$|\langle z^{*},0-0\rangle-\langle x^{*},x^{\prime}-x\rangle-\langle y^{*},\sigma(x^{\prime})-\sigma(x)\rangle|\leq\epsilon\|z^{*},x^{*},y^{*}\|\|(x^{\prime}-x,\sigma(x^{\prime})-\sigma(x),0-0)\|$
according to the semismoothness property of $F$ at $(x,\sigma(x),0)$. By possibly decreasing $\delta$ we may also assume that for all $x^{\prime}\in\tilde{U}$ with $(x^{\prime},\sigma(x^{\prime}))\in{\cal B}_{\delta}(x,\sigma(x))$ and all $(g_{x}^{T},g_{y}^{T})\in\Phi(x^{\prime},\sigma(x^{\prime}))$ there holds
$|\varphi(x^{\prime},\sigma(x^{\prime}))-\varphi(x,\sigma(x))-\langle g_{x},x^{\prime}-x\rangle-\langle g_{y},\sigma(x^{\prime})-\sigma(x)\rangle|\leq\epsilon\|(x^{\prime}-x,\sigma(x^{\prime})-\sigma(x))\|.$
Choose $\delta^{\prime}\in(0,\delta/\sqrt{1+\gamma^{2}})$ such that ${\cal B}_{\delta^{\prime}}(x)\subset U$ and consider $x^{\prime}\in{\cal B}_{\delta^{\prime}}$ and ${u^{*}}^{T}\in\Psi(x^{\prime})$ together with $(g_{x}^{T},g_{y}^{T})\in\Phi(x^{\prime},\sigma(x^{\prime}))$ and $x^{*}\in\Sigma(x^{\prime},\sigma(x^{\prime}))(g_{y})$ such that $u^{*}=g_{x}^{T}+{x^{*}}^{T}$. Further choose $z^{*}$ such that $(z^{*},x^{*},-g_{y})\in\mathrm{gph}\,D^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ according to the definition of $\Sigma$. Then $(x^{\prime},\sigma(x^{\prime})\in{\cal B}_{\delta}(x,\sigma(x))$ and consequently
$|\varphi(x^{\prime},\sigma(x^{\prime}))-\varphi(x,\sigma(x))-\langle u^{*},x^{\prime}-x\rangle|$
$\leq|\varphi(x^{\prime},\sigma(x^{\prime}))-\varphi(x,\sigma(x))-\langle g_{x},x^{\prime}-x\rangle-\langle g_{y},\sigma(x^{\prime})-\sigma(x)\rangle|+|\langle g_{y},\sigma(x^{\prime})-\sigma(x)\rangle-\langle x^{*},x^{\prime}-x\rangle|$
$\leq\epsilon\|(x^{\prime}-x,\sigma(x^{\prime})-\sigma(x))\|+\epsilon\|(z^{*},x^{*},-g_{y})\|\|(x^{\prime}-x,\sigma(x^{\prime})-\sigma(x),0-0)\|$
$\leq\epsilon\|x^{\prime}-x\|\sqrt{1+\gamma^{2}}(1+\sqrt{1+\gamma^{2}}\|g_{y}\|)\leq\epsilon\|x^{\prime}-x\|\sqrt{1+\gamma^{2}}(1+\sqrt{1+\gamma^{2}}M),$
verifying that $\theta$ is semismooth with respect to $\Psi$ at $x$. This completes the proof. ∎
Remark 5.4.
In the setting of Theorem 5.3, consider a point $x\in U$ where $F$ is semismooth at $(x,\sigma(x),0)$ and consider the linear functions $\varphi_{i}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$, $\varphi_{i}(x,y)=y_{i}$, $i=1,\ldots,m$, together with mappings $\tilde{\Psi}_{i}:U\rightrightarrows\mathbb{R}^{1\times n}$ given by $\tilde{\Psi}_{i}(u):=\{{u^{*}}^{T}\,\big{|}\,u^{*}\in\Sigma(u,\sigma(u))(-e_{i})$, where $e_{i}$ denotes the $i$-th unit vector in $\mathbb{R}^{m}$. It follows from Theorem 5.3 that $\sigma_{i}(u)=\varphi_{i}(u,\sigma(u))$ is semismooth at $x$ with respect to $\tilde{\Psi}_{i}$, $i=1,\ldots,m$. Hence $\sigma$ is also semismooth with respect to the mapping $\tilde{\Psi}$, which assigns to each $u\in U$ the set of matrices $A$ whose $i$-th row is an element from $\tilde{\Psi}_{i}(u)$.
Corollary 5.5.
In the setting of Theorem 5.1, assume that we are given an open set $\Omega$ and a continuous mapping $\sigma:\Omega\to\mathbb{R}^{m}$ such that for every $x\in\Omega$ the mapping $F$ is semismooth at $(x,\sigma(x),0)$, the qualification condition
$(x^{*},0)\in D^{*}F(x,\sigma(x),0)(z^{*})\ \Rightarrow z^{*}=0,x^{*}=0$
holds and there exists a neighborhood $U\times V$ of $(x,\sigma(x))$ such that $\mathrm{gph}\,S\cap(U\times V)=\mathrm{gph}\,\sigma\cap(U\times V)$. Then the following statements hold:
  1. 1.
    $\sigma$ is locally Lipschitz on $\Omega$.
  2. 2.
    $\sigma$ is semismooth on $\Omega$ with respect to some mapping $\tilde{\Psi}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n\times m}$, which can be constructed analogously to Remark 5.4.
  3. 3.
    Consider a function $\varphi:D\to\mathbb{R}$, where $D$ is an open subset of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ containing $\mathrm{gph}\,\sigma$ and assume that $\varphi$ is semismooth on $\mathrm{gph}\,\sigma$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$. Then the mapping $\theta:\Omega\to\mathbb{R}$ given by $\theta(x)=\varphi(x,\sigma(x))$ is semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ given by
    $\Psi(x):=\begin{cases}\{g_{x}^{T}+{x^{*}}^{T}\,\big{|}\,x^{*}\in\Sigma(x,\sigma(x))(g_{y}),(g_{x}^{T},g_{y}^{T})\in\Phi(x,\sigma(x))\}&\mbox{if $x\in\Omega$,}\\ \emptyset&\mbox{else.}\end{cases}$
Statement 3 enables us to apply the ImP approach to the numerical solution of a class of problems (2) with nonsmooth objectives $\varphi$ via a suitable bundle method. Apart from the case when $\sigma$ is a single-valued and Lipschitzian localization of $S$ for all $x\in\Omega$, however, the construction of $\tilde{\Psi}$ according to Remark 5.4 is rather complicated. We will now show that the theory considerably simplifies in the case of SCD mappings.
For an SCD mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, the subspaces $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ belong to ${\cal Z}_{m,n+m}$ and therefore $L^{*}\subset\mathbb{R}^{m+n+m}$. By identifying $\mathbb{R}^{m+n+m}$ with $\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$, we denote the elements of $L^{*}$ by $(z^{*},x^{*},y^{*})\in\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$. We start our analysis with the following result:
Proposition 5.6.
For every subspace $L^{*}\in{\cal Z}_{m,{n+m}}$ satisfying
$(z^{*},x^{*},0)\in L^{*}\Rightarrow z^{*}=0,\ x^{*}=0$ (13)
there are unique matrices $Z_{L^{*}}\in\mathbb{R}^{m\times m}$ and $X_{L^{*}}\in\mathbb{R}^{n\times m}$ such that
$L^{*}={\rm rge\;}(Z_{L^{*}},X_{L^{*}},I)=\{(Z_{L^{*}}p,X_{L^{*}}p,p)\,\big{|}\,p\in\mathbb{R}^{m}\}.$ (14)
In addition, for the norm
$\kappa(L^{*}):=\big{\|}\begin{pmatrix}Z_{L}^{*}\\ X_{L^{*}}\end{pmatrix}\big{\|}$
there holds
$\kappa(L^{*})=\sup\{\|(z^{*},x^{*})\|\,\big{|}\,(z^{*},x^{*},y^{*})\in L^{*},\|y^{*}\|\leq 1\}.$
Proof.
Consider an $(m+n+m)\times m$ matrix $B$ whose columns form a basis for the $m$-dimensional subspace $L^{*}\subset R^{m+n+m}$. We can partition $B$ into the matrices $B_{z^{*}}\in\mathbb{R}^{m\times m}$, $B_{x^{*}}\in\mathbb{R}^{n\times m}$ and $B_{y^{*}}\in\mathbb{R}^{m\times m}$ so that
$B=\begin{pmatrix}B_{z^{*}}\\ B_{x^{*}}\\ B_{y^{*}}\end{pmatrix}.$
We claim that condition (13) ensures the nonsingularity of the matrix $B_{y^{*}}$. Assume on the contrary that there is some $0\not=p\in\mathbb{R}^{m}$ satisfying $B_{y^{*}}p=0$. Then we have $Bp=(B_{z^{*}}p,B_{x^{*}}p,0)\in L^{*}$ and consequently $B_{z^{*}}p=0$, $B_{x^{*}}p=0$ due to (13). Thus $Bp=0$, which is not possible since the colums of $B$ are a basis for $L^{*}$ and are therefore linearly independent. Hence the matrix $B_{y^{*}}$ is nonsingular and the columns of the matrix $BB_{y^{*}}^{-1}$ form another basis for $L^{*}$. Representation (14) now follows with
$Z_{L^{*}}=B_{z^{*}}B_{y^{*}}^{-1},\ X_{L^{*}}=B_{x^{*}}B_{y^{*}}^{-1}$
and it is easy to see that it is unique. Finally,
$\sup\{\|(z^{*},x^{*})\|\,\big{|}\,(z^{*},x^{*},y^{*})\in L^{*},\|y^{*}\|\leq 1\}=\sup\{\|(Z_{L^{*}}y^{*},X_{L^{*}}y^{*})\|\,\big{|}\,\|y^{*}\|\leq 1\}=\big{\|}\begin{pmatrix}Z_{L}^{*}\\ X_{L^{*}}\end{pmatrix}\big{\|}$
and we are done. ∎
Definition 5.7.
We denote by ${\cal Z}_{m,n+m}^{\rm reg}$ the collection of all subspaces $L^{*}\in{\cal Z}_{m,n+m}$ satisfying (13).
Remark 5.8.
From (14) it easily follows that for any $L^{*}\in{\cal Z}_{m,n+m}^{\mathrm{reg}\,}$ the corresponding primal subspace $L\in{\cal Z}_{n+m,m}$ has the representation
$L=\{(p_{x},X_{L^{*}}^{T}p_{x}+Z_{L^{*}}^{T}p_{z},p_{z})\,\big{|}\,(p_{x},p_{z})\in\mathbb{R}^{n}\times\mathbb{R}^{m}\}.$
In the following statements we collect some useful relationships concerning the subspaces $L^{*}$ from ${\cal Z}_{m,n+m}^{\mathrm{reg}\,}$ and the moduli $\kappa(L^{*})$ and ${\rm scd\,reg\,}F(x,y,z)$ (introduced below).
Lemma 5.9.
Assume that $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ has the SCD property at $(x,y,z)\in\mathrm{gph}\,F$. Then the following statements are equivalent:
  1. (i)
    $(z^{*},x^{*},0)\in(\bigcup{\mathcal{S}}^{*}F)(x,y,z)\Rightarrow z^{*}=0,\ x^{*}=0.$ (15)
  2. (ii)
    ${\mathcal{S}}^{*}F(x,y,z)\subset{\cal Z}_{m,n+m}^{\rm reg}$.
  3. (iii)
    ${\rm scd\,reg\,}F(x,y,z):=\sup\{\|(z^{*},x^{*})\|\,\big{|}\,(z^{*},x^{*},y^{*})\in(\bigcup{\mathcal{S}}^{*}F)(x,y,z),\|y^{*}\|\leq 1\}<\infty.$
In this case one also has
${\rm scd\,reg\,}F(x,y,z):=\sup\{\kappa(L^{*})\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,y,z)\}.$ (16)
Proof.
The implication (i)$\Rightarrow$(ii) follows immediately from Proposition 5.6. We prove the implication (ii)$\Rightarrow$(iii). Assuming that (iii) does not hold there are sequences $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in L_{k}^{*}\subset(\bigcup{\mathcal{S}}^{*}F)(x,y,z)$ verifying $\|y_{k}^{*}\|\leq 1$ for all $k$ and $\|(z_{k}^{*},x_{k}^{*})\|\to\infty$ as $k\to\infty$. Since $L^{*}_{k}$ is a subspace, we have $(\tilde{z}^{*}_{k},\tilde{x}^{*}_{k},\tilde{y}^{*}_{k}):=(z_{k}^{*},x_{k}^{*},y_{k}^{*})/\|(z_{k}^{*},x_{k}^{*},y_{k}^{*})\|\in L_{k}^{*}$ for all $k$. Taking into account that the metric space ${\cal Z}_{m,n+m}$ is compact, by possibly passing to a subsequence we can assume that $L_{k}^{*}$ converges to some $L^{*}\in{\cal Z}_{m,n+m}$ and $(\tilde{z}^{*}_{k},\tilde{x}_{*}^{k},\tilde{y}^{*}_{k})$ converges to some $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\in L$ with $\|(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\|=1$. Since ${\mathcal{S}}^{*}F(x,y,z)$ is closed in ${\cal Z}_{m,n+m}$, there holds $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and because of $\tilde{y}_{k}^{*}/\|z_{k}^{*},x_{k}^{*},y_{k}^{*}\|\to 0$ we have $\tilde{y}^{*}=0$. Hence $(\tilde{z}^{*},\tilde{x}^{*},0)\in L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$ and $\|\tilde{z}^{*},\tilde{x}^{*})\|=1$ showing $L^{*}\not\in{\cal Z}_{m,n+m}^{\rm reg}$. This proves the implication (ii)$\Rightarrow$(iii). The implication (iii)$\Rightarrow$(i) holds trivially. Finally, formula (16) is a consequence of the definitions of ${\rm scd\,reg}F$, $(\bigcup{\mathcal{S}}^{*}F)$ and $\kappa(L^{*})$. ∎
Proposition 5.10.
Assume that $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ has the SCD property around some point $(x,y,z)\in\mathrm{gph}\,F$. Then
$\limsup_{(x^{\prime},y^{\prime},z^{\prime}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,F}}(x,y,z)}{\rm scd\,reg\,}F(x^{\prime},y^{\prime},z^{\prime})\leq{\rm scd\,reg\,}F(x,y,z).$ (17)
Proof.
If $\gamma:={\rm scd\,reg\,}F(x,y,z)=\infty$ there is nothing to prove. Thus we can suppose that $\gamma<\infty$ and consequently, by Lemma 5.9, condition (15) is fulfilled. Assume now on the contrary that the inequality (17) does not hold and we can find sequences $(x_{k},y_{k},z_{k})\to(x,y,z)$ and $(z_{k}^{*},x_{k}^{*},y_{k}^{*})\in L_{k}^{*}\in{\mathcal{S}}^{*}F(x_{k},y_{k},z_{k})$ with $\|y_{k}\|^{*}\leq 1$ and $\limsup_{k\to\infty}\|(z^{*}_{k},x^{*}_{k})\|=:\bar{\gamma}>\gamma$. If $\bar{\gamma}$ is finite, by possibly passing to a subsequence we can assume that $(z_{k}^{*},x_{k}^{*},y_{k}^{*})$ converges to some $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})$ with $\bar{\gamma}=\|(\tilde{z}^{*},\tilde{x}^{*})\|$. By possibly passing to a subsequence once more we can assume that $L_{k}^{*}$ converges to some $L^{*}\in{\mathcal{S}}^{*}F(x,y,z)$. Since $(\tilde{z}^{*},\tilde{x}^{*},\tilde{y}^{*})\in L^{*}$ and $\|\tilde{y}^{*}\|\leq 1$, we obtain the contradiction $\gamma<\bar{\gamma}=\|(\tilde{z}^{*},\tilde{x}^{*})\|\leq\kappa(L^{*})\leq\gamma$. If $\bar{\gamma}=\infty$ we can employ the same arguments already used in the proof of Lemma 5.9 to obtain a contradiction. ∎
The next statement pertains to the $\Psi$-ss property of a single-valued continuous selection of (31) in the SCD case.
Theorem 5.11.
Consider a mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ and a point $(\bar{x},\bar{y})$ with $0\in F(\bar{x},\bar{y})$. Assume that $F$ has the SCD property around $(\bar{x},\bar{y},0)$ and ${\rm scd\,reg\,}F(\bar{x},\bar{y},0)<\infty$. Further suppose that there exists a neighborhood $\Omega$ of $\bar{x}$ and a continuous function $\sigma:\Omega\to\mathbb{R}^{m}$ verifying $\sigma(\bar{x})=\bar{y}$ and $0\in F(x,\sigma(x))$, $x\in\Omega$. Then there exists an open neighborhood $U\subset\Omega$ of $\bar{x}$ such that for every $x\in U$ there holds ${\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$ and $\sigma$ is semismooth at $x$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$ given by
$\Psi(x):=\begin{cases}\{-X_{L^{*}}^{T}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\}&\mbox{if $x\in U$},\\ \emptyset&\mbox{else,}\end{cases}$
provided $F$ is SCD semismooth at $(x,\sigma(x),0)$.
Proof.
Due to the definition of the SCD property around a point and Proposition 5.10 we can find an open neighborhood $U^{\prime}\times V^{\prime}\times W^{\prime}$ of $(\bar{x},\bar{y},0)$ such that for every $(x,y,z)\in U^{\prime}\times V^{\prime}\times W^{\prime}$ the mapping $F$ has the SCD property around $(x,y,z)$ and ${\rm scd\,reg\,}F(x,y,z)<\infty$. Next consider the open neighborhood $U^{\prime\prime}:=\{x\in\Omega\cap U^{\prime}\,\big{|}\,\sigma(x)\in V^{\prime}\}$ and choose an open neighborhood $U$ of $\bar{x}$ satisfying $\mathrm{cl}\,U\subset U^{\prime\prime}$. This choice of $U$ ensures that ${\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$, $x\in U$. Next we show that $\Psi$ is locally bounded at $x\in U$. Indeed, we have $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)<\infty$ and, due to the continuity of $\sigma$ and Proposition 5.10, we can find a neighborhood $\tilde{U}\subset\Omega$ of $x$ satisfying ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in\tilde{U}$. Hence, $\|X^{*}\|<\gamma+1$ for every $X^{*}\in\Psi(x^{\prime})$ and our claim is established. In order to show the last assertion, consider $x\in U$ such that $F$ is SCD semismooth at $(x,\sigma(x),0)$, set $\gamma:={\rm scd\,reg\,}F(x,\sigma(x),0)$ and choose $\tilde{\delta}>0$ such that ${\rm scd\,reg\,}F(x^{\prime},\sigma(x^{\prime}),0)<\gamma+1$ $\forall x^{\prime}\in{\cal B}_{\tilde{\delta}}(x)$. By definition of the SCD semismooth property we can find for every $\epsilon>0$ some $\delta\in(0,\tilde{\delta})$ such that for every $x^{\prime}\in{\cal B}_{\delta}(x)$, every subspace $L^{*}\in{\mathcal{S}}^{*}F(x^{\prime},\sigma(x^{\prime}),0)$ and every $p\in\mathbb{R}^{m}$ we have
$|\langle Z_{L^{*}}p,0-0\rangle-\langle X_{L^{*}}p,x^{\prime}-x\rangle-\langle p,\sigma(x^{\prime})-\sigma(x)\rangle|$ $=|\langle p,X_{L^{*}}^{T}(x^{\prime}-x)+\sigma(x^{\prime})-\sigma(x)\rangle|$
$\leq\epsilon\|(x^{\prime}-x,\sigma(x^{\prime})-\sigma(x),0-0)\|\|(Z_{L^{*}}p,X_{L^{*}}p,p)\|$
$\leq\epsilon(\|x^{\prime}-x\|+\|\sigma(x^{\prime})-\sigma(x)\|)\sqrt{1+\kappa(L^{*})^{2}}\|p\|$
and $\kappa(L^{*})\leq\gamma+1$, showing
$\|X_{L^{*}}^{T}(x^{\prime}-x)+\sigma(x^{\prime})-\sigma(x)\|\leq\epsilon(\|x^{\prime}-x\|+\|\sigma(x^{\prime})-\sigma(x)\|)\gamma^{\prime}$
with $\gamma^{\prime}=\sqrt{1+(1+\gamma)^{2}}$. Thus $(1-\epsilon\gamma^{\prime})\|\sigma(x^{\prime})-\sigma(x)\|\leq\|x^{\prime}-x\|(\|X_{L^{*}}^{T}\|+\epsilon\gamma^{\prime})$ implying $\|\sigma(x^{\prime})-\sigma(x)\|\leq 2\|x^{\prime}-x\|((\gamma+1)+\frac{1}{2})$ whenever $\epsilon\in(0,1/(2\gamma^{\prime}))$. Hence
$\|X_{L^{*}}^{T}(x^{\prime}-x)+\sigma(x^{\prime})-\sigma(x)\|\leq\epsilon\|x^{\prime}-x\|\gamma^{\prime}(2\gamma+4),$
verifying the $\Psi$-ss property of $\sigma$ at $x$. This completes the proof. ∎
Corollary 5.12.
Consider a mapping $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ and a continuous mapping $\sigma:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, satisfying $0\in F(x,\sigma(x))$, $x\in\Omega$. Further assume that for every $x\in\Omega$ the mapping $F$ is SCD semismooth at $(x,\sigma(x),0)$ and $\mathrm{scd\,reg}\,F(x,\sigma(x),0)<\infty$. Then $\sigma$ is semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n\times m}$ given by
$\Psi(x):=\begin{cases}\{-X_{L^{*}}^{T}\,\big{|}\,L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\}&\mbox{if $x\in\Omega$},\\ \emptyset&\mbox{else.}\end{cases}$
In particular, for every function $\varphi:D\to\mathbb{R}$ defined on an open set $D\subset\mathbb{R}^{n}\times\mathbb{R}^{m}$ containing $\mathrm{gph}\,\sigma$, which is semismooth on $\mathrm{gph}\,\sigma$ with respect to some mapping $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{1\times(n+m)}$, there holds that the function $\theta:\Omega\to\mathbb{R}$, $\theta(x):=\varphi(x,\sigma(x))$ is semismooth on $\Omega$ with respect to the mapping $\Theta:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$ given by
$\Theta(x):=\begin{cases}\{g_{x}^{T}-g_{y}^{T}X_{L^{*}}^{T}\,\big{|}\,(g_{x}^{T},g_{y}^{T})\in\Phi(x,\sigma(x)),L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)\}&\mbox{if $x\in\Omega$},\\ \emptyset&\mbox{else.}\end{cases}$
Note that in Corollary 5.5 we require that $\sigma$ is some continuous localization of the solution map $S$, whereas in Corollary 5.12 the mapping $\sigma$ only needs to be any continuous selection of $S$. This is illustrated by the following academic example.
Example 5.13.
Consider the mapping $F:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ given by $F(x,y)=|y|-|x|$. Then the solution map $S$ to the equation $F(x,y)=0$ is given by
$S(x)=\{x\}\cup\{-x\}$
and no single-valued localization of $S$ around the origin exists. On the other hand, consider a continuous selection $\sigma$, e.g., $\sigma(x)=\{x\}$. It is easy to compute by means of Lemma 2.13 that
${\mathcal{S}}^{*}F(x,\sigma(x),0)=\begin{cases}\{{\rm rge\;}\big{(}1,(-1,1)\big{)}\}&\mbox{if $x\not=0$,}\\ \{{\rm rge\;}\big{(}1,(-1,1)\big{)},{\rm rge\;}\big{(}1,(1,1)\big{)},{\rm rge\;}\big{(}1,(-1,-1)\big{)},{\rm rge\;}\big{(}1,(1,-1)\big{)}\}&\mbox{if $x=0$.}\end{cases}$
We see that all subspaces $L^{*}\in{\mathcal{S}}^{*}F(x,\sigma(x),0)$, $x\in\mathbb{R}$ belong to ${\cal Z}_{1,2}^{\rm reg}$ and therefore $\sigma$ is semismooth on $\mathbb{R}$ with respect to the mapping
$\Psi(x)=\begin{cases}\{1\}&\mbox{if $x\not=0$,}\\ \{1,-1\}&\mbox{if $x=0$.}\end{cases}$
We see that $\nabla\sigma(x)\subset\Psi(x)$, $x\in\mathbb{R}$, but at the origin $\Psi(0)$ also contains the value $-1$ which has nothing to do with the derivative of $\sigma$. Note that we can arbitrarily modify $\Psi$ on a finite subset of $\mathbb{R}$ without destroying the semismooth property.
6 Conclusion
The paper contains a deep analysis of some generalizations of the classical semismoothness property. Namely, it deals with the so-called $\Psi$-semismoothness and new notions of $\Gamma$-semismoothness and SCD-semismoothness. These notions are closely related to each other and have a number of useful properties concerning, in particular, some fundamental numerical approaches to nonsmooth problems. For instance, whereas the standard rules of generalized differential calculus attain typically the form of inclusions, the presence of the mentioned semismoothness properties enables us to construct suitable approximations of the considered mappings which fulfill the respective calculus rules as equalities. This is especially helpful in connection with a family of bundle methods in nonsmooth optimization.
A special attention has been paid to the semismoothness of localizations (selections) of solution maps to parameter-dependent inclusions and their possible compositions with real-valued objectives. The presented analysis of these mappings paves the way to an efficient usage of the ImP approach in a broad class of problems with equilibrium constraints.